### v1

# Dynamics of Two Species with Density-Dependent Interactions in a Mutualistic Context

Abstract
Mutualistic interactions, where individuals from different species can benefit from each other, are widespread across ecosystems. This study develops a general deterministic model of mutualism involving two populations, assuming that mutualism may involve both costs and benefits for the interacting individuals, leading to density-dependent effects on the dynamics of the two species. This framework aims at generalizing pre-existing models, by allowing the ecological interactions to transition from mutualistic to parasitic when the respective densities of interacting species change. Through ordinary differential equations and phase portrait analysis, we derive general principles governing these systems, identifying sufficient conditions for the emergence of certain dynamic behaviors. In particular, we show that limit cycles can arise when interactions include parasitic phases but are absent in strictly mutualistic regimes. This framework provides a general approach for characterizing the population dynamics of interacting species and highlights the effect of the transitions from mutualism to parasitism due to density dependence.
*Corresponding author. E-mail: chloe.mian@univ-grenoble-alpes.fr
Keywords. Parasitism, Limit cycles, Phase portrait analysis, Population interactions
Mathematics Subject Classification. 37N25, 92-10, 34C23, 92D25
1 Introduction
In natural communities, interactions between individuals from different species are widespread, ranging from antagonistic relationships, such as competition or parasitism, to cooperative ones, such as mutualism (Bronstein,, 2015; Janzen,, 1985). Mutualistic interactions include processes such as seed dispersal (Janzen,, 1985), pollination (Klein et al.,, 2007), resource exchange (van der Heijden et al.,, 2008), and protection (Bronstein,, 2015), where individuals from both species may benefit from the relationship. For example, mycorrhizal fungi provide plants with essential nutrients, while plants provide carbohydrates to the interacting fungi (Bonfante and Genre,, 2010). Similarly, in pollination, plants produce nectar feeding some insects, which in turn transport pollen and facilitate plant reproduction (Sargent and Ackerly,, 2008).
Mutualistic interactions can be either facultative or obligatory. If facultative, the dynamics of each species can benefit from each other, but each species can be maintained in the absence of the other (e.g. generalist pollination, where plants and pollinators can often find alternative partners (Sargent and Ackerly,, 2008)). Otherwise, if mutualistic interactions are obligatory, the persistence of each species entirely depends on the mutualistic relationship and hence on the presence of the interacting species (e.g. the fig-wasp relationship, Bronstein, 2001).
The nature and strength of the interactions can also depend on population densities. Variations in species densities can shift the cost-benefit dynamics of the relationship, consequently shifting interactions along a continuum from mutualism to parasitism. For example, in ant-plant mutualism, plants provide ants with food in exchange for protection. However, at low ant densities, plants may not receive sufficient protection, while at high densities, the cost of supporting the ants may outweigh the protective benefits (Bronstein,, 1994; Holland et al.,, 2002; Janzen,, 1985). Similarly, in pollination systems, the investment of plants in nectar can become inefficient if pollinator densities are too low to ensure effective pollination, or too high, leading to increased competition among pollinators and diminishing returns for the plants (Kawai and Tokeshi,, 2007; Morris et al.,, 2010; Holland et al.,, 2002).
Mathematical models of ecological interactions have mostly focused on antagonistic relationships such as predator-prey and competitive dynamics. While some models addressed mutualism, they often assumed strictly positive interactions, leading to unrealistic outcomes like unbounded population growth or infinite densities (Brauer and Castillo-Chavez,, 2012; Hale and Valdovinos,, 2021; May,, 1972). Indeed, because of the complexity and context-dependence of mutualistic interactions, modeling them as a direct extension of Lotka–Volterra systems overlooks key mechanisms, such as density-dependent saturation, that are essential to preventing unbounded growth. Moreover, the factors shaping the persistence of mutualistic interactions and the transitions between parasitic and mutualistic interactions driven by populations densities variations (i.e., on ecological time scales) are much less documented than in predator–prey systems (Holland et al.,, 2002; Revilla and Encinas-Viso,, 2015). Density dependence and saturation effects to prevent unbounded growth or systematic extinction were considered in Brauer and Castillo-Chavez, (2012), bounded growth terms were added in Holland et al., (2002) to model that mutualistic benefits naturally diminish as population sizes increase, and finally Ruiz-Herrera, (2025) analyzed mutualistic systems with a density-dependent framework, emphasizing the role of benefit–cost ratios in shaping long-term interaction outcomes. Asymmetries in obligate vs. facultative mutualistic interactions among species have also been shown to have a significant influence on system stability and resilience (Bronstein,, 1994; Holland and DeAngelis,, 2010; Neuhauser and Fargione,, 2004). Most of these previous models studying the dynamics of mutualistic interactions rely on ordinary differential equations to describe the conditions for coexistence and stability (see Hale and Valdovinos, (2021) for a review). Other approaches such as network analysis have also been used to explore interspecific relationships within ecological systems (Bascompte et al.,, 2003) and partial differential equations to model the spatial dynamics of mutualistic interactions (Aliyu and Mohd,, 2021).
Examples of mutualistic interactions with density-dependent effect have been documented in natural populations (Bronstein,, 2015). Yet, no general theory has been developed to study the coexistence of mutualistic species while accounting for this important process. As pointed out by Hale and Valdovinos, (2021), it is striking that all mutualistic interactions models studied so far show similar outcomes and behaviors, suggesting that some general properties can be found, independently of the peculiarities of each model. For example, there must be some general conditions for the stable coexistence of the two species when the interaction is density-dependent. Previous models of mutualistic interactions often focus on equilibrium points, typically identifying stable nodes or saddle points as the primary outcomes of these systems with particular modeling choices of the functional forms relating the rate of interactions, the birth and death rates, and the populations densities. While these models provide valuable insights, they rarely consider the possibility of more complex dynamics, such as limit cycles or oscillatory behaviors, therefore preventing general predictions.
By proposing and analyzing a general mathematical model of two-species mutualism, we aim at tackling these limitations, focusing in particular on two main questions:
  1. 1.
    What types of dynamical behaviors, including equilibrium configurations, can be derived from models of mutualistic interactions, and which types are not covered by these models?
  2. 2.
    To what extent does incorporating transitions between mutualism and parasitism allow us to observe previously neglected dynamics?
To address these questions, Section 2 first provides the general definitions of density-dependent effects acting on the dynamics of species involved in ecological interactions, which will be used throughout the paper. We thus introduce a system of differential equations for two interacting populations, incorporating density-dependent interactions, without focusing on specific types of interactions within a particular ecological context and on limited range of parameters. In Section 3, we formulate general and robust assumptions derived from the analysis of deterministic mutualistic models (Theorem 3.1). We then present examples of mutualistic interaction models that, despite aiming to represent different aspects of mutualism or emphasizing certain features over others, fall within the framework of Theorem 3.1. In Section 4, we use the extended definition of mutualistic interactions and identify sufficient conditions for the emergence of oscillatory dynamics. These conditions, often overlooked in previous studies, provide new insights into the potential for stable coexistence or extinction. After discussing our results in Section 5, we present the corresponding proofs in Section 6.
2 Model And Key Definitions
2.1 General Framework
We model the dynamics of two interacting populations using the system:
$\left\{\begin{array}[]{ll}\dot{x}&=xf\left(x,y\right)\\ \dot{y}&=yg\left(x,y\right)\end{array}\right$. (1)
where $x$ and $y$ represent the densities of each species, $f$ and $g$ describe the density-dependent growth rates of each species. Because we model population sizes, the analysis is confined to the non-negative quadrant of $\mathbb{R}^{2}$. This form is commonly used to describe two-species dynamics without accounting for immigration, as in the models of Hale and Valdovinos, (2021), May, (1972), and Brauer and Castillo-Chavez, (2012).
2.2 Key Assumptions and Definition
To analyze and interpret the behavior of the system, we follow an approach similar to that used in Brauer and Castillo-Chavez, (2012), relying on the partial derivatives of the growth functions. These derivatives capture how the growth rate of each species responds to changes in the density of both species. For this reason, starting from Section 3.2, we will assume that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$ functions i.e. continuous and continuously differentiable. We focus on specific regions of the phase plane defined by the population densities $x$ and $y$. In each region, the sign of the partial derivatives of the interaction terms determines the local effect of one species on the growth rate of the other. As a result, an interaction may be positive in one part of the phase portrait and negative in another, depending on the local population densities. Different regions of the phase portrait may correspond to different interaction regimes, described below:
  • Mutualism and Parasitism: $\frac{\partial f}{\partial y}>0$ and $\frac{\partial g}{\partial x}>0$ represent a region of strict mutualism, where each species positively influences the growth of the other. We refer to the model as strictly mutualistic when the condition is satisfied for all population densities $x$ and $y$. Conversely, $\frac{\partial f}{\partial y}<0$ or $\frac{\partial g}{\partial x}<0$ indicates a region of parasitism, as the presence of one species negatively impacts the growth of the other.
  • Intraspecific Competition: $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$ describe negative feedback within each species, representing competition, where individuals of the same population limit each other’s growth.
  • Intraspecific Cooperation: $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ describe positive feedback within each species, representing cooperative interactions, where individuals of the same population support each other’s growth.
  • Dynamic Transitions: The signs of $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ are not fixed and may change with species densities, reflecting a continuum of interactions ranging from mutualism to parasitism. The phase portrait may contain regions where the signs of $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ differ, reflecting changes in interaction type.
In contrast to many pre-existing models where these partial derivatives always remain within the same strict sign or even remain constant regardless of species densities (Hale and Valdovinos,, 2021), we allow the sign of these partial derivatives to change. Such assumption allows for transitions from mutualistic (positive value) to parasitic interactions (negative value) between species, depending on their respective densities. We thus give an extended definition of mutualism as follows:
Definition 2.1.
A system of differential equations of the form (1) will be said to be mutualistic if, in the phase portrait of $\mathbb{R}_{+}\times\mathbb{R}_{+}$, there is at least a region where $\frac{\partial g}{\partial x}>0$ and at least a region where $\frac{\partial f}{\partial y}>0$. These two regions may be disjoint.
These definitions are illustrated in Figure 1, which presents two different situations, both consistent with mutualism as defined in Definition 2.1. In the first case, a strictly mutualistic region appears, although the condition does not hold for all densities $x$ and $y$. In the second case, the system remains mutualistic overall, but no strictly mutualistic region exists; that is, both populations benefit from the interaction, though not simultaneously.
(a) Local region of strict mutualism: Overlapping positive partial derivatives.
(b) Disjoint regions of mutualistic influence: Non-overlapping mutualistic effects.
Figure 1: Illustration of density-dependent effects of mutualism (Definition 2.1) through the signs of cross partial derivatives. In both situations, the interaction is classified as mutualistic according to Definition 2.1, even though the local effects of one species on the other vary. Figure 1(a): $\frac{\partial f}{\partial y}>0$ everywhere, meaning that species $y$ always has a positive effect on the growth of species $x$. The hatched region indicates where $\frac{\partial g}{\partial x}>0$, so that species $x$ also benefits species $y$ locally. This region represents a strictly mutualistic zone. Outside this zone, the effect of species $x$ on $y$ is negative, yet the system still qualifies as mutualistic under Definition 2.1. Figure 1(b): The cross partial derivatives $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ vary and are positive in distinct, non-overlapping regions: the hatched (resp. dashed) region corresponds to $\frac{\partial g}{\partial x}>0$ (resp. $\frac{\partial f}{\partial y}>0$). Outside these regions, the interaction is locally parasitic.
3 General dynamics of two species with any types of interactions
In this work, we study the behavior of system (1), including stability, oscillations, and other dynamics, with a particular focus on how the specific forms of $f$ and $g$ influence species coexistence. In Section 3.1 we consider a general framework in which $f$ and $g$ are assumed to be continuous only, and we apply index theory (see Perko, (2013) and Appendix A for details) to prove Theorem 3.1. Then, using our general Theorem 3.1, we study the different assumptions described in Section 2.2 adding the assumption that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$ to investigate more precisely how the nature of the interactions considered can affect the outcomes of the system. We mainly focus on simple equilibrium points and use methods based on the Jacobian matrix to analyze their stability and nature. For example, regardless of the specific nature of the interaction represented by system (1), certain information can already be deduced about the equilibrium point $\left(0,0\right)$ by applying these methods (see Appendix B and Brauer and Castillo-Chavez, (2012)).
3.1 Assumptions and Equilibrium Index Alternation
We introduce several assumptions about the functions $f$ and $g$ that are commonly derived from biological models of population dynamics. Most of these assumptions, inspired by models reviewed in Hale and Valdovinos, (2021), are robust and remain valid even when $f$ and $g$ are subject to minor local perturbations. We recall that the zeros of $f$ and $g$ define the isoclines of the system, and their intersections determine the equilibrium points.
Hypothesis H1.
The functions $f$ and $g$ each define a single additional isocline, beyond those at $x=0$ and $y=0$: a vertical isocline for $f$ and a horizontal isocline for $g$. The zeros $\left(x,y\right)$ of $f$ (resp. of $g$) form a single continuous curve.
We will denote these curves by $\Gamma_{f}$ and $\Gamma_{g}$:
$\Gamma_{f}:=\{\left(x,y\right)\in\mathbb{R}_{+}\times\mathbb{R}_{+},f\left(x,y\right)=0\}$
and
$\Gamma_{g}:=\{\left(x,y\right)\in\mathbb{R}_{+}\times\mathbb{R}_{+},g\left(x,y\right)=0\},$
with $\Gamma_{f}\neq\emptyset$ and $\Gamma_{g}\neq\emptyset$.
Hypothesis H2.
$\Gamma_{f}$ and $\Gamma_{g}$ delimit regions of strict constant sign of $f$ and $g$. Moreover we impose that the signs of $f$ (respectively of $g$) change on either side of the curve $\Gamma_{f}$ (respectively of the curve $\Gamma_{g}$).
Hypothesis H3.
No more than two isoclines intersect at any given point.
In particular, H3 implies that $\left(0,0\right)\notin\Gamma_{f}\cup\Gamma_{g}$.
Hypothesis H4.
We exclude the boundary cases where the equilibrium points are formed by two isoclines that only touch at that point but do not cross, and the case where the isoclines are coincident.
These assumptions directly influence the properties of the vector field. One characteristic of equilibrium points is their index, which describes the local behavior of the flow in their neighborhood. The index helps determine whether an equilibrium point acts as an attractive, a repulsive, or a saddle point, and reveals how trajectories interact with the point (see Perko, (2013) and Appendix A for more details). In particular, for systems satisfying H1 to H4 on $f$ and $g$, the indices of equilibrium points in the positive quadrant exhibit a regular alternation along the isoclines. This result is formalized in the following theorem:
Theorem 3.1.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying hypotheses H1 to H4. Then, in the positive quadrant, the equilibrium points of the system alternate along the isoclines between having an index of +1+1 and an index of 1-1.
The proof of this result, provided in Section 6, is illustrated in Figure 2.
(a) Local structure of the phase portrait around an index +1+1 equilibrium.
(b) Local structure of the phase portrait around an index 1-1 equilibrium.
Figure 2: Two possible local configurations of the phase portrait near an equilibrium point, corresponding to the two admissible clockwise sign changes of the vector field across adjacent regions. Illustration of Proof of Theorem 3.1. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. Note: A single arrow is used to represent the direction of the vector field within each region for simplicity. For instance, in the $\left(+,+\right)$ region, the direction of the vector field $\left(\nearrow\right)$ may vary from horizontal $\left(\rightarrow\right)$ (excluded) to vertical $\left(\uparrow\right)$ (excluded), as long as the signs remain strictly positive.
Remark 3.1.
Here, we recall that we did not assume the functions $f$ and $g$ to be $\operatorname{\mathcal{C}^{1}}$; we only required the continuity of the curves $\Gamma_{f}$ and $\Gamma_{g}$. For instance, we may consider a function $f$ in which the birth rate of species $x$ depends on the density of species $y$ through a threshold-like mechanism—linear at low densities of $y$, and saturating to a constant value at high densities.
Remark 3.2.
We recall that the model cover the wide parasitic–mutualistic continuum. The interactions may therefore represent competition, predation, or mutualism, as long as the functions $f$ and $g$ satisfy the conditions stated in assumptions H1H4.
Additional assumptions on $f$ and $g$ are necessary to completely describe the configuration and outcomes of the systems. In fact, knowing that certain points have an index +1+1 is not enough to determine whether an equilibrium is attractive, repulsive, or even a center. Moreover, identifying the nature of an equilibrium point with index +1+1 is not sufficient to determine the nature of all other equilibrium points with the same index. Without further assumptions on $f$ and $g$, two equilibrium points with index +1+1 may exhibit different behaviors, with one being attractive and the other repulsive.
To illustrate these findings, we present an example of a system that can model a mutualistic interaction, and shows an alternation of equilibrium points with different characteristics.
Example 3.1.
Let us consider the following system:
$\left\{\begin{array}[]{ll}\dot{x}&=xf\left(x,y\right)=x\left(e_{1}y-d_{1}\left(\left(x-a_{1}\right)^{b_{1}}+c_{1}\right)\right),\\ \dot{y}&=yg\left(x,y\right)=y\left(e_{2}x-d_{2}\left(\left(y-a_{2}\right)^{b_{2}}+c_{2}\right)\right),\end{array}\right$. (2)
with $(a_{i},c_{i},d_{i},e_{i},i\in\{1,2\})$ being positive constants, and $(b_{i},i\in\{1,2\})$ being positive integers.
The functions $f$ and $g$ satisfy assumptions H1 to H4. We choose $\frac{\partial f}{\partial y}=e_{1}>0$ and $\frac{\partial g}{\partial x}=e_{2}>0$ to represent strict mutualism with a constant positive effect. The other parameters represent internal limitations $d_{i}$, constraints related to critical thresholds $a_{i}$, or fixed costs $c_{i}$. We choose the $b_{i}$ to be even so that the partial derivative $\frac{\partial f}{\partial x}=-d_{1}b_{1}(x-a_{1})^{b_{1}-1}$ can change sign, in this example, depending on the position of $x$ relative to the threshold $a_{1}$.
Figure 3 displays the phase portrait corresponding to (2), for certain parameter values. In the positive quadrant, there is an alternation of equilibrium points with index +1+1 and 1-1, with +1+1 points of different natures, including a repulsive point and an attractive point. This system could model mutualistic interactions where there is stable coexistence of the two species if the initial population densities are sufficiently high. However, if the density of one species is too low, the populations collapse, the state $(0,0)$ being attractive.
Figure 3: Example of a mutualism model exhibiting different types of equilibrium points. Phase portrait of system (2) for the parameter values: $a_{1}=6$, $b_{1}=2$, $c_{1}=1$, $d_{1}=1$, $e_{1}=1$ and $a_{2}=5$, $b_{2}=4$, $c_{2}=1$, $d_{2}=1/4$, $e_{2}=1$. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. The figure illustrates a mutualistic system with multiple types of equilibrium points.
3.2 Intraspecific Competition and Cooperation Assumption
From this section, we assume that the functions $f$ and $g$ are of class $\operatorname{\mathcal{C}^{1}}$. By applying the implicit function theorem and using the fact that the relevant partial derivatives have strict signs locally, we obtain information about the local monotonicity of the $x$-isocline $\Gamma_{f}$ and the $y$-isocline $\Gamma_{g}$. These results are summarized in Table 1.
In many models presented in Hale and Valdovinos, (2021), intraspecific competition is a key mechanism in regulating population dynamics within each species, preventing the uncontrolled population growth which may be criticized in some models of mutualistic interactions. With this additional assumption, all previous equilibrium points with an index of +1+1 become attractive points.
Corollary 3.1.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ both $\operatorname{\mathcal{C}^{1}}$ satisfying hypotheses H1 to H4, and $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$. Then, in the positive quadrant, the system exhibits an alternation of attractive nodes and saddle points along the isoclines.
Remark 3.3.
For a model satisfying the conditions of Corollary 3.1, if there are at least two equilibrium points in the positive quadrant, then an attractive point exists. This implies the existence of conditions under which coexistence of the two species is possible.
Under the opposite constant sign assumptions, i.e. intraspecific cooperation for both species as described in 2.2, the system in the positive quadrant alternates between repulsive nodes and saddle points along the isoclines.
Corollary 3.2.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ both $\operatorname{\mathcal{C}^{1}}$ satisfying hypotheses H1 to H4, and $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$. Then, in the positive quadrant, the system exhibits an alternation of repulsive nodes and saddle points along the isoclines.
This reflects a dynamic where intraspecific cooperation leads to instability in some regions, while stability is limited to saddle point behavior in others. A saddle point represents an unstable equilibrium where the system converges along one direction (stable) but diverges along another (unstable). Under these assumptions, this reflects a fragile coexistence scenario where small perturbations can push the system away from equilibrium, leading to instability in the dynamics of the two species.
Thus, negative assumptions on the sign of the partial derivatives $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$ (which biologically correspond to intraspecific competition, always negative regardless of species densities) prevent the emergence of any coexistence solutions other than attractive equilibrium points or saddle points.
3.3 Illustrative Examples of Theorem 3.1
In this section, we illustrate Theorem 3.1 using four models of mutualistic interactions reviewed in Hale and Valdovinos, (2021), each represented on Figure 4 (in the order presented), each emphasizing a different characteristic of mutualism. These four models use different mathematical representations to emphasize certain aspects of mutualism over others, and describe facultative (Neuhauser and Fargione,, 2004) or obligate (Graves et al.,, 2006), density-dependent, or saturating mutualistic interactions. Although their mathematical formulation reflects different perspectives on mutualistic interactions, as discussed in Section 2.2, we show that they satisfy the assumptions of system (1), highlighting the generality of the framework provided in our study.
  • (a)
    In Zhang, (2003) the author models the dynamics of two mutualistic species using the same mathematical formulation twice, reflecting an interaction between species at the same trophic level. In this model, the growth rate of each species is influenced by a quadratic term that depends on the density of the other species. The intraspecific competition is represented by a constant term, while the interaction between species varies in sign: it remains positive up to a threshold, beyond which it becomes negative. This formulation captures the effect of saturation and illustrates the continuum between mutualism and parasitism. The system is given by:
    $\begin{cases}\dot{x}=r_{x}x\left(c_{x}-x-a_{x}(y-b_{x})^{2}\right),\\ \dot{y}=r_{y}y\left(c_{y}-y-a_{y}(x-b_{y})^{2}\right),\end{cases}$ (3)
    where $r_{x}$, $r_{y}$, $a_{x}$, $a_{y}$, $c_{x}$, $c_{y}$ are positive real constants and $b_{x}$, $b_{y}$ are real constants. Here, there is the assumption of intraspecific competition, where $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$. Moreover, the interaction is mutualistic as long as $y$ remains below the threshold $b_{x}$, but when $y$ exceeds this value, $\frac{\partial f}{\partial y}$ turns negative. This dynamic highlights how the strength and nature of the interaction shift depending on species densities.
  • (b)
    Neuhauser and Fargione, (2004) introduce a model describing facultative mutualism between plants ($x$) and mycorrhizal fungi ($y$). In this formulation, the effects on species $y$ remain constant, regardless $x$ or $y$, with intraspecific competition always negative ($\frac{\partial g}{\partial y}<0$) and the mutualistic benefit from $x$ always positive ($\frac{\partial g}{\partial x}>0$). Similarly, intraspecific competition for $x$ is constant and negative ($\frac{\partial f}{\partial x}<0$). However, the term $\frac{\partial f}{\partial y}$, which characterizes the nature of the interaction, varies depending on the densities of both species. Specifically, as $y$ increases, the benefit for $x$ diminishes and can eventually turn negative, shifting the interaction away from mutualism. However, this effect is modulated by the presence of $x$: for huge values of $x$, $\frac{\partial f}{\partial y}$ is more strongly positive, reinforcing mutualistic benefits. $x$ and $y$ must increase together to prevent the relationship from becoming parasitic. The system is given by:
    $\begin{cases}\dot{x}=r_{x}x\left(\frac{K_{x}\gamma_{xy}y-x}{K_{x}+\gamma_{xy}y}-ay\right),\\ \dot{y}=r_{y}y\left(\frac{K_{y}+\alpha_{yx}x-y}{K_{y}}\right),\end{cases}$ (4)
    where all parameters are positive real constants.
  • (c)
    The model proposed by Graves et al., (2006) represents obligate mutualism for lichens, where the benefits of the interaction are tightly linked to the densities of both species. The system assumes constant negative intraspecific competition, with identical functional forms for both species, and the mutualistic interaction varies with the density of the other species. Specifically, as the density of $y$ increases, the effect on $x$ decreases, but without any saturation effect—just a continuous increase with density.
    The system is given by:
    $\begin{cases}\dot{x}=x\left(r_{x0}+(r_{x1}-r_{x0})\left(1-\exp{(-k_{x}y)}\right)-a_{x}x\right),\\ \dot{y}=y\left(r_{y0}+(r_{y1}-r_{y0})\left(1-\exp{(-k_{y}x)}\right)-a_{y}y\right).\end{cases}$ (5)
    where $r_{x1}$, $r_{y1}$, $a_{x}$, $a_{y}$, $k_{x}$, $k_{y}$ are positive real constants and $r_{x0}\leq r_{x1}$, $r_{y0}\leq r_{y1}$, $r_{x0}$ and $r_{y0}$ are real constants. Here, the interaction benefits are regulated by an exponential term, which prevents unbounded growth as species densities increase.
  • (d)
    A more general consumer-resource approach is presented by Holland and DeAngelis, (2010), which models the bidirectional exchange of benefits in plant-mycorrhizal systems. In this framework, one species provides resources while the other enhances nutrient uptake efficiency. The system follows a similar functional form for both species, but the dynamics are more complex as all aspects—including the intraspecific competition—vary with the densities of both species.
    The system is given by:
    $\begin{cases}\dot{x}=x\left(r_{x}+c_{x}\left(\frac{a_{xy}y}{h_{y}+y}\right)-q_{x}\left(\frac{\beta_{xy}y}{e_{x}+x}\right)-s_{x}x\right),\\ \dot{y}=y\left(r_{y}+c_{y}\left(\frac{a_{yx}x}{h_{x}+x}\right)-q_{y}\left(\frac{\beta_{yx}x}{e_{y}+y}\right)-s_{y}y\right).\end{cases}$ (6)
    where all parameters are positive constants, except $r_{x}$ and $r_{y}$, which are real constants. In this model, the mutualistic interaction is obligate when the intrinsic growth rates are negative ($r_{x}\leq 0$, $r_{y}\leq 0$). The interaction structure is dynamic, with the intraspecific competition for $x$ decreasing as $y$ grows, but increasing as $x$ increases. Conversely, the effect of $y$ on $x$ increases with $x$, but decreases as $y$ increases. This creates a more flexible framework for describing how species’ interactions evolve with changing densities.
Depending on the biological aspects these models aim to represent, the mathematical forms differ, leading to a varying number of equilibrium points. However, all of them satisfy conditions H1 to H4 and exhibit an alternation of equilibrium point indices along the isoclines, as a result leading to one or more stable coexistence configurations.
(a) Mutualism at the same trophic level.
(b) Plant–mycorrhiza, facultative mutualism.
(c) Obligate mutualism in lichens.
(d) Consumer-resource approach.
Figure 4: Phase portraits of four differents models of mutualism. Each model highlights different aspects of mutualism while satisfying the assumptions H1 to H4 of system (1) and exhibiting an alternation of equilibrium point indices along the isoclines. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. (4(a)) Parameters for Zhang, (2003) model, in (3): $c_{x}=7$, $c_{y}=8$, $a_{x}=0.7$, $a_{y}=0.5$, $b_{x}=4.5$, $b_{y}=3$, $r_{x}=7$, $r_{y}=4$. (4(b)) For Neuhauser and Fargione, (2004) model, in (4): $r_{x}=1.4$, $r_{y}=1$, $K_{x}=2$, $K_{y}=5$, $\gamma_{xy}=4$, $\alpha_{xy}=1.5$, $a=0.2$. (4(c)) For Graves et al., (2006) model, in (5): $r_{x0}=-2$, $r_{y0}=0.5$, $r_{x1}=4$, $r_{y1}=5$, $k_{x}=0.2$, $k_{y}=0.2$, $a_{x}=0.35$, $a_{y}=0.5$. (4(d)) For Holland and DeAngelis, (2010) model, in (6) : $r_{x}=2$, $r_{y}=3$, $c_{x}=3$, $c_{y}=1$, $a_{xy}=1$, $a_{yx}=2.6$, $h_{x}=1$, $h_{y}=1$, $q_{x}=1$, $q_{y}=2.7$, $\beta_{xy}=1$, $\beta_{yx}=1.5$, $e_{x}=0.5$, $e_{y}=3$, $s_{x}=0.7$, $s_{y}=0.5$.
4 Periodic solutions in the dynamics of two mutualistic species
A system of the form (1) can represent either predator-prey, competition, or mutualistic interactions. The results described so far do not distinguish between these different types of interactions.
In this section, we clarify the role of extended mutualism (Definition 2.1) and contrast it with the more constrained definition of strict mutualism. Specifically, we demonstrate that under the strict mutualism assumptions, where both $\frac{\partial f}{\partial y}>0$ and $\frac{\partial g}{\partial x}>0$ in a same region of the positive quadrant, the system cannot give rise to periodic solutions such as limit cycles in its neighborhood.
Remark 4.1.
Here we aim at investigating attractor limit cycles. A limit cycle is an isolated closed trajectory that can be stable, unstable or half-stable for nearby trajectories. When the limit cycle is stable, trajectories that start close to it tend to converge towards it over time, which stabilizes the oscillation. Thus, attractive limit cycles represent stable periodic oscillations, where populations repeatedly return to the same cycle of variation. In mutualistic interactions, the presence of limit cycles might suggest that the populations of both species oscillate over time in a stable way, without extinction or explosive growth. Hereafter, we will refer to limit cycles exclusively as attractive limit cycles.
Remark 4.2.
Limit cycles differ from centers, such as those observed in the Lotka-Volterra predator-prey system. Centers are equilibrium points where the trajectories of the system form closed, periodic orbits, but they are neither attractive nor repulsive. As a result, the behavior of the system is highly sensitive to initial conditions: each new set of initial conditions produces a distinct trajectory, making the system highly sensitive to small variations (May,, 1972). However, even in the case of centers, when the Lotka-Volterra predator-prey model is modified to represent mutualistic interactions, the centers disappear.
We recall that periodic solutions can only surround a region with an index of +1+1 (Section 3.1, Figure 2(a)). However, equilibrium points do not necessarily generate an attractive limit cycle around them. Limit cycle will arise around a repulsive equilibrium point, where nearby trajectories diverge from the equilibrium, but are eventually captured by the limit cycle.
4.1 Case of strict mutualism
We first study whether a limit cycle can be observed under the assumption that the repulsive equilibrium point is strictly mutualistic, that is, under the assumption $\frac{\partial f}{\partial y}>0$ and $\frac{\partial g}{\partial x}>0$.
Theorem 4.1.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying hypotheses H1 to H4. We assume that the curves $\Gamma_{f}$ and $\Gamma_{g}$ intersect at a repulsive equilibrium point $\left(x^{*},y^{*}\right)$, with $\frac{\partial f}{\partial y}(x^{*},y^{*})>0$ and $\frac{\partial g}{\partial x}(x^{*},y^{*})>0$. Then, there exists a region $R_{1}$ in the phase portrait containing this point, in which the partial derivatives do not change sign, and no limit cycle exists that surrounds this point within $R_{1}$.
A similar result can be found with an attractive equilibrium point surrounded by a repulsive limit cycle. If a repulsive limit cycle exist, then an attractive limit cycle could be found around this point, with an alternation between an attractive point, a repulsive limit cycle, and then an attractive limit cycle.
Theorem 4.2.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying hypotheses H1 to H4. We assume that the curves $\Gamma_{f}$ and $\Gamma_{g}$ intersect at an attractive equilibrium point $\left(x^{*},y^{*}\right)$, with $\frac{\partial f}{\partial y}(x^{*},y^{*})>0$ and $\frac{\partial g}{\partial x}(x^{*},y^{*})>0$. Then, there exists a region $R_{2}$ in the phase portrait containing this point, in which the partial derivatives do not change sign, and no limit cycle exists that surrounds this point within $R_{2}$.
Thus, around a point with index +1 (attractive or repulsive), in a system (1) where the conditions on the partial derivatives do not change, we cannot find periodic solutions such as limit cycles. However, the Definition 2.1 of extended mutualism includes the case of strict mutualism. We simply deduce that in the region where both species have a strictly positive effect on each other and in which the partial derivatives do not change sign, no periodic solutions such as limit cycles can arise. We will now explore what happens outside this region.
4.2 Case of extended mutualism
In this section, we use the extended definition of mutualism provided in Definition 2.1, which includes the previous case of strict mutualism and describes partially mutualistic interactions, where the interactions between species can be mutualistic only for certain combinations of species densities. As previously discussed, the emergence of a limit cycle can indeed only occur in a context where the interaction is either parasitic or competitive, and not in a region where both $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ simultaneously and remain unchanged. Here, we consider density-dependent effects that give rise to regions where the mutualistic interaction locally shifts to a parasitic one. We identify two different scenarios where the nature of the interaction is not fixed, i.e., where $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ change signs, leading to the emergence of a limit cycle.
We also introduce the assumption that intraspecific competition can change sign and become intraspecific cooperation (for example, $\frac{\partial g}{\partial y}$ may change sign). In mutualism, this can occur if an increase in the population of species $x$ leads to a higher demand for individuals of species $y$ to support the interaction. To investigate this scenario, we focus on a single equilibrium point in the population quadrant, excluding the one at $(0,0)$ and those on the axes.
For our first configuration, we make the following assumptions:
  1. 1.1
    $\frac{\partial f}{\partial x}<0$: Intraspecific competition, the density of species $x$ is always limited by the intraspecific competition among individuals.
  2. 1.2
    $\frac{\partial g}{\partial y}>0$ then $\frac{\partial g}{\partial y}<0$: At low density, species $y$ experiences self-cooperation rather than competition, but at higher density, negative effects do emerge.
  3. 1.3
    $\frac{\partial g}{\partial x}>0$: Species $x$ always has a positive effect on the density of species $y$.
  4. 1.4
    $\frac{\partial f}{\partial y}>0$ then $\frac{\partial f}{\partial y}<0$: At low density, species $y$ has a positive effect on species $x$, but as its density increases, the interaction shifts from mutualism to parasitism.
  5. 1.5
    $f\left(0,y_{1}\right)=0$: Above the threshold $y_{1}$, the population of $x$ declines, regardless of whether its own density is low or high.
  6. 1.6
    $g\left(x_{1},0\right)=0$: The threshold $x_{1}$ of species $x$ required for species $y$ to persist at low density.
  7. 1.7
    $f\left(x_{2},0\right)=0$: There exists an equilibrium at $x_{2}$ where species $x$ can persist in the absence of species $y$. Beyond $x_{2}$, the species $x$ declines due to overpopulation.
  8. 1.8
    $x_{1}<x_{2}$: The threshold density for species $x$ to persist in isolation ($x_{2}$) is higher than the threshold where species $x$ can sustain $y$ ($x_{1}$). Otherwise, species $y$ goes extinct.
  9. 1.9
    $\left(x^{*},y^{*}\right)$ is a repulsive equilibrium point:
    $x^{*}\frac{\partial f}{\partial x}\left(x^{*},y^{*}\right)+y^{*}\frac{\partial g}{\partial y}\left(x^{*},y^{*}\right)>0,$
    and
    $x^{*}y^{*}\left(\frac{\partial f}{\partial x}\left(x^{*},y^{*}\right)\frac{\partial g}{\partial y}\left(x^{*},y^{*}\right)-\frac{\partial f}{\partial y}\left(x^{*},y^{*}\right)\frac{\partial g}{\partial x}\left(x^{*},y^{*}\right)\right)>0.$
    This condition ensures that small perturbations around $(x^{*},y^{*})$ will lead to divergence, favoring oscillatory or cyclic behavior.
  10. 1.10
    The isoclines $f=0$ and $g=0$ have the shapes illustrated in Figure 5.
    Figure 5: Phase portrait assuming extended mutualism leading to a cyclic behaviour (illustration of 1.10 of the rectangular region $ABCD$). The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. By assumption 1.9, $(x^{*},y^{*})$ is a repulsive point (the white point), which implies, by Theorem 3.1, that $(x_{2},0)$ is a saddle point (the grey points), as well as $(0,0)$.
Remark 4.3.
These assumptions are more restrictive than those in Section 3.1 and imply assumptions H1 to H4.
Under conditions 1.1-1.10, we have the following result, which will be proven in Section 6.
Theorem 4.3.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 1.1-1.10. Then a limit cycle exists inside the positive quadrant.
Example 4.1.
Let us consider the following system:
$\left\{\begin{array}[]{ll}\dot{x}&=xf\left(x,y\right)=x\left(a_{1}-b_{1}\left(y-c_{1}\right)^{2}-d_{1}x\right),\\ \dot{y}&=yg\left(x,y\right)=y\left(-a_{2}-b_{2}\left(y-c_{2}\right)^{2}+d_{2}x\right),\end{array}\right$. (7)
with $(a_{i},b_{i},c_{i},d_{i},i\in\{1,2\})$ being positive constants. For a certain range of parameters, the functions $f$ and $g$ will satisfy assumptions 1.1 to 1.10. However, according to assumptions 1.9 and 1.10, $\Gamma_{f}$ and $\Gamma_{g}$ must intersect while both being decreasing in order to obtain a repulsive equilibrium point. Lowering the threshold $c_{2}$, increasing the threshold $c_{1}$, or decreasing the interaction strength $b_{1}$ leads to a shift in the equilibrium location, making it attractive instead. System (7) is illustrated in Figure 6.
Figure 6: Phase portrait leading to a limit cycle. Simulation of system (7) with parameters: $a_{1}=3.7$, $b_{1}=0.5$, $c_{1}=2$, $d_{1}=0.6$ and $a_{2}=3$, $b_{2}=0.2$, $c_{2}=4$, $d_{2}=1$. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. The white point represents a repulsive point and gray points represent saddle points. The purple trajectory starts at a population density of 2 for species $x$ and 6 for species $y$. It converges toward the limit cycle from the outside, moving away from the extinction states, either of both species or the extinction of $y$. The gray trajectory starts at densities of 7 for species $x$ and 10 for species $y$, and converges toward the limit cycle from the inside, moving away from the interior equilibrium point.
For our second configuration, we make the following assumptions:
  1. 2.1
    $\frac{\partial g}{\partial y}<0$: Intraspecific competition for species $y$.
  2. 2.2
    $\frac{\partial f}{\partial x}>0$ then $\frac{\partial f}{\partial x}<0$: At low density, species $x$ experiences self-cooperation rather than competition, but at higher density, negative effects emerge.
  3. 2.3
    $\frac{\partial f}{\partial y}>0$: Species $y$ always has a positive effect on species $x$.
  4. 2.4
    $\frac{\partial g}{\partial x}<0$ then $\frac{\partial g}{\partial x}>0$: At low density, species $x$ has a negative effect on species $y$, but as its density increases, the interaction shifts from parasitism to mutualism.
  5. 2.5
    $f(0,y_{1})=0$: Beyond the threshold $y_{1}$, the population of $x$ grows due to mutualism, regardless of its own density.
  6. 2.6
    $g\left(0,y_{2}\right)=0$: There exists an equilibrium at $y_{2}$ where species $y$ can persist in the absence of species $x$. Beyond $y_{2}$, the species $y$ declines due to overpopulation.
  7. 2.7
    $y_{1}<y_{2}$: The threshold density for species $y$ to persist in isolation ($y_{2}$) is higher than the threshold where species $y$ can sustain $x$ ($y_{1}$). Otherwise, species $x$ goes extinct.
  8. 2.8
    $\left(x^{*},y^{*}\right)$ is a repulsive equilibrium point.
  9. 2.9
    The isoclines $f=0$ and $g=0$ have the shapes illustrated in Figure 7.
    Figure 7: Phase portrait assuming partial mutualism leading to a cyclic behaviour (illustration of 2.9 of the rectangular region $ABCD$)). By assumption 2.8, $(x^{*},y^{*})$ is a repulsive point (white point), which implies that $(0,0)$ and $(0,y_{2})$ are saddle points (grey points). The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange.
Remark 4.4.
As before, these assumptions are more restrictive than those presented in Section 3.1 and imply assumptions H1 to H4.
Under conditions 2.1-2.10, we have the following result, which will be proven in Section 6.
Theorem 4.4.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 2.1-2.10. Then a cycle limit exists inside the positive quadrant.
4.3 From Parasitism to Mutualism: Parameter-Driven Transitions in Interaction Dynamics
In this section, we include models that, while not always strictly mutualistic, satisfy Definition 2.1, and fit to our general framework.
Depending on the position in the phase portrait and the population densities of each species, a model satisfying these assumptions could represent a parasitic interaction in one region, followed by a competitive interaction in another region and with different species densities, then a mutualistic one, and then return to a parasitic interaction at yet other densities. As Theorem 3.1 is stated in a more general context, it remains valid despite the specific transition dynamics examined in Section 2.2. Additionally, Corollaries 3.1 and 3.2 characterize the nature of equilibrium points with index +1+1 when intraspecific competition remains constant in sign. However, a given model representing a specific type of interaction can describe an entirely different interaction when its parameter values change (as for example intraspecific competition) while still remaining within the framework of Theorem 3.1.
In this section, we examine a first model which, depending on the signs of its coefficients—and thus its partial derivatives—can represent either a parasitic or a mutualistic interaction, as discussed in Section 2.2. We analyze the consequences of this change on phase portraits and equilibrium points of index +1+1. We then consider a second model that falls within Definition 2.1, where we keep the coefficient signs unchanged but slightly modify the values of certain parameters. We observe that an equilibrium previously identified as attractive in a parasitic region remains attractive but now in a mutualistic region for both species. The following examples are consistent with assumptions H1 to H4.
Our first example is the following system:
$\begin{cases}\dot{x}=\delta x\left(a_{x}-b_{x}\left(x-c_{x}\right)^{2}-d_{x}y\right),\\ \dot{y}=\delta y\left(a_{y}-b_{y}\left(y-c_{y}\right)^{4}-d_{y}x\right),\end{cases}$ (8)
where:
$\delta=\begin{cases}1&\text{for parasitism}\\ -1&\text{for mutualism}.\end{cases}$
Depending on the sign of $\delta$, the same model can represent either a parasitic interaction, where $\frac{\partial g}{\partial x}<0$ and $\frac{\partial f}{\partial y}<0$, or a mutualistic interaction, where both $\frac{\partial g}{\partial x}>0$ and $\frac{\partial f}{\partial y}>0$.
Figure 8 presents two phase portraits of this model, illustrating the effect of changing the sign of $\delta$. This modification transforms the system from a parasitic interaction—where a stable equilibrium allows both species to coexist, two additional equilibria enable either species to outcompete the other and persist alone, and extinction at $(0,0)$ is an attractor—to a mutualistic interaction. In the mutualistic case, a stable coexistence equilibrium remains, but at lower population densities. Extinction becomes repulsive, and if initial densities are sufficiently high, population sizes can grow explosively.
While the alternation predicted by the theorem remains unchanged, the nature of equilibrium points of index +1+1 shifts: previously attractive points become repulsive, and vice versa.
(a) Phase portrait of parasitism of (8), $\delta=1$.
(b) Phase portrait of mutualism of (8), $\delta=-1$.
Figure 8: Phase portraits of system (8) illustrating how changes in parameter signs alter the interaction type. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. With: $a_{x}=8$, $b_{x}=1$, $c_{x}=5.5$, $d_{x}=1$ and $a_{y}=9.5$, $b_{y}=0.5$, $c_{y}=4$, $d_{y}=1$.
In our second example, we do not change the sign of the partial derivatives; the model has two regions, a mutualistic region and a parasitic region. We maintain the nature of the interaction but slightly modify the values of certain constants, which shifts the positions of the equilibrium points. We use the following model:
$\begin{cases}\dot{x}=x(-a_{x}-b_{x}(x-c_{x})^{2}+d_{x}y),\\ \dot{y}=y(a_{y}-b_{y}(x-c_{y})^{4}-d_{y}y).\end{cases}$ (9)
Figure 9 provides two phase portraits of the same model, demonstrating how varying parameter values affect equilibrium locations. The growth region of $\Gamma_{g}$ corresponds to $\frac{\partial g}{\partial x}>0$, while its decline region corresponds to $\frac{\partial g}{\partial x}<0$, with $\frac{\partial g}{\partial y}<0$ always. A symmetric structure applies to $f$.
Thus, in Figure 9(a), the stable equilibrium lies in a region where the interaction is parasitic for species $y$ and beneficial for species $x$. In contrast, in Figure 9(b), stable coexistence occurs within a region of strict mutualism.
(a) Phase portrait of (9), with parameters: $a_{x}=1.7$, $c_{x}=5.2$ and $c_{y}=5$.
(b) Phase portrait of (9), with parameters: $a_{x}=1.5$, $c_{x}=1.9$ and $c_{y}=7$.
Figure 9: Phase portraits illustrating how changing parameter values—while maintaining the interaction type—modifies the equilibrium locations. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. Black points represent attractive points and gray points represent saddle points. System (9) with: $b_{x}=0.4$, $d_{x}=1$, $a_{y}=6.5$, $b_{y}=0.03$ and $d_{y}=1$.
5 Discussion
In this study we investigated the deterministic dynamics of two interacting species under density-dependent interactions, with a particular focus on mutualistic interactions. Our approach allowed to draw a general theoretical framework, formalized by Theorem 3.1, which is independent of the specific type of ecological interaction considered, covering for instance predator-prey, competitive and mutualistic interactions. However, while most mutualistic models reviewed in Hale and Valdovinos, (2021) satisfy assumptions H1 to H4, some mutualistic models do not necessarily adhere to these assumptions. For instance, they may involve more than two isoclines intersecting at a single point, or include emigration effects (Thompson et al.,, 2006; Holland and DeAngelis,, 2010). Our framework nevertheless provides general predictions of the dynamics of species involved in a wide array of ecological interactions, and highlights the key-role of density-dependent processes, that are likely widespread in natural communities (Kawatsu and Kondoh,, 2018).
We extended the classical definition of mutualism, allowing density-dependent transitions between mutualistic and antagonistic interactions. We thus proposed a broader Definition 2.1: a system is considered mutualistic if there exists at least one region for each species in the phase portrait where each benefits from the presence of the other. This relaxed definition enables the exploration of more diverse ecological dynamics and highlights biologically realistic features of species interactions. In textbook example of mutualistic interactions, such as pollination of flowering plants by insects, large densities of invasive bees have been documented to decrease plant reproductive success because of the damages caused by over-visitation of flowers (Aizen et al.,, 2014). As shown in some of the mutualistic models we studied, it is possible that small changes in parameter values could cause an equilibrium point that was initially located in a strictly mutualistic region to shift into a parasitic one. By allowing such transitions between mutualism and parasitism, our extended framework opened to the analysis of more complex systems and revealed new dynamical phenomena such as the emergence of limit cycles. Our theoretical predictions therefore provide a more realistic view of population dynamics of interacting species, opening new research avenues for community ecology.
6 Proofs
In this section, we provide the proofs of the different results stated in Section 3 and 4. We begin with the proof of Theorem 3.1, which describes the alternation of equilibrium point indices along the isoclines.
Proof of Theorem 3.1.
Let us consider the dynamical system (1), where $f$ and $g$ satisfy hypotheses H1 to H4.
From Assumption H1, the strictly positive quadrant contains exactly two continuous curves $\Gamma_{f}$ and $\Gamma_{g}$, which may intersect. These curves partition $\mathbb{R}_{+}\times\mathbb{R}_{+}$ into distinct regions. According to H2, inside each zone, the signs of $f$ and $g$, therefore those of $\dot{x}$ and $\dot{y}$ (for $x>0$ and $y>0$), remain constant. Thus, each region in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ can be associated with a unique pair of signs $\left(\dot{x},\dot{y}\right)$. There are four possible pairs of signs: $\left(+,+\right)$, $\left(-,-\right)$, $\left(+,-\right)$ and $\left(-,+\right)$ for $\left(\dot{x}>0,\dot{y}>0\right)$, $\left(\dot{x}<0,\dot{y}<0\right)$, $\left(\dot{x}>0,\dot{y}<0\right)$ and $\left(\dot{x}<0,\dot{y}>0\right)$.
By H3, equilibrium points correspond to the intersections of $\Gamma_{f}$ and $\Gamma_{g}$. Since there are no more than two curves intersecting at any given point, the plane near an equilibrium point is divided into four regions. To move from one zone to another, crossing a $\Gamma_{f}$ curve causes a change in the sign of $\dot{x}$, crossing a $\Gamma_{g}$ curve causes a change in the sign of $\dot{y}$.
The assumption that the sign of either $\dot{x}$ or $\dot{y}$ changes on either side of an isocline (H2) ensures that at least one component of the pair $\left(\dot{x},\dot{y}\right)$ changes sign. By Assumption H4, since the isoclines do not coincide, both components of $\left(\dot{x},\dot{y}\right)$ cannot change sign at the same time. From these properties, near an equilibrium point, there are exactly two possible clockwise sequences of sign changes:
  • (i)
    $\left(+,+\right)$, $\left(+,-\right)$, $\left(-,-\right)$ and $\left(-,+\right)$,
  • (ii)
    $\left(+,+\right)$, $\left(-,+\right)$, $\left(-,-\right)$ and $\left(+,-\right)$.
Each pair corresponds to a specific direction of the vector field. We can thus associate four phase portraits with each sign alternation, representing the rotation of sign pairs near the equilibrium point in Figure 2.
Using this information, we can determine the index of the vector field around the equilibrium point. Given the form of the vector field in Figure 2(b), an equilibrium point with the sign alternation pattern (ii) corresponds to an index of 1-1, indicating a saddle point. Similarly, according to the form of the vector field in Figure 2(a), an equilibrium point with the sign alternation pattern (i) corresponds to an index of +1+1, which can represent an attractive point, a repulsive point, or even a center.
Finally, H1 ensures that only two isoclines, $\Gamma_{f}$ and $\Gamma_{g}$, exist in the strictly positive quadrant, making equilibrium points their intersections. H4 requires that the curves alternate at each intersection due to their crossing. Consequently, equilibrium points with an index of +1+1 are always followed by points with an index of 1-1 along the isoclines, and vice versa. ∎
We now provide the proof of Corollary 3.1, which is a consequence of Theorem 3.1. The proof of Corollary 3.2 is very similar and will be omitted.
Proof of Corollary 3.1.
Let $\left(x,y\right)$ an equilibrium point in the strictly positive quadrant. We can calculate the Jacobian of the system (1) at the point $\left(x,y\right)$:
$J\left(x,y\right)=\begin{pmatrix}x\frac{\partial f\left(x,y\right)}{\partial x}&x\frac{\partial f\left(x,y\right)}{\partial y}\\ y\frac{\partial g\left(x,y\right)}{\partial x}&y\frac{\partial g\left(x,y\right)}{\partial y}\end{pmatrix}$. (10)
And the trace of $J\left(x,y\right)$ is:
$\text{Tr}\left(J\left(x,y\right)\right)=x\frac{\partial f\left(x,y\right)}{\partial x}+y\frac{\partial g\left(x,y\right)}{\partial y}<0.$ (11)
It implies that the point $\left(x,y\right)$ is an attractive node if $\text{det}\left(J\left(x,y\right)\right)>0$, else, if $\text{det}\left(J\left(x,y\right)\right)<0$, the point $\left(x,y\right)$ is a saddle point. ∎
We now prove why stable limit cycles cannot be observed under assumptions of strict mutualism.
Proof of Theorem 4.1.
$\left(x^{*},y^{*}\right)$ is a repulsive equilibrium point with the assumptions $\frac{\partial f}{\partial y}(x^{*},y^{*})>0$ and $\frac{\partial g}{\partial x}(x^{*},y^{*})>0$. According to Table 1 (see Appendix C), this case only arises when, in addition, $\frac{\partial f}{\partial x}(x^{*},y^{*})>0$ and $\frac{\partial g}{\partial y}(x^{*},y^{*})>0$. Since the functions $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$, there exists a neighborhood $R_{1}$ around the point $(x^{*},y^{*})$ such that $\frac{\partial f}{\partial y}>0$, $\frac{\partial g}{\partial x}>0$, $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$.
By the implicit function theorem, we conclude that, in the neighborhood of the point $(x^{*},y^{*})$, both isoclines $\Gamma_{f}$ and $\Gamma_{g}$ are decreasing. To correspond to a repulsive equilibrium point (with index +1+1), the sign alternation around this point must match that of Figure 2(a). There are only two possible phase portrait configurations in the neighborhood see Figure 10.
(a) First possible alternation
(b) Second possible alternation
Figure 10: Illustration of Proof of Theorem 4.1. Two possible neighborhoods of the repulsive equilibrium point $(x^{*},y^{*})$ that could lead to the formation of a limit cycle around it.
Let us consider the case of the Figure 10(a). The other case of the Figure 10(b) follows similarly with $g$. We know the sign of $f$ and $g$ on either side of $\Gamma_{f}$ and $\Gamma_{g}$. Let $u$ and $v$ be points as shown in Figure 10(a), with respective coordinates $\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix}$ and $\begin{pmatrix}x_{2}\\ y_{2}\end{pmatrix}$, where $x_{1}<x_{2}$ and $y_{1}<y_{2}$, and define the function $h$ as follows:
$h:[0,1]$ $\to\mathbb{R}$ (12)
$t$ $\mapsto f(u+t(v-u))$ (13)
Then, we have:
$h(0)$ $=f(u)>0,$ (14)
$h(1)$ $=f(v)<0.$ (15)
By construction, $h$ is continuously differentiable $\operatorname{\mathcal{C}^{1}}$, with derivative given by:
$h^{\prime}(t)=(x_{2}-x_{1})\frac{\partial f}{\partial x}(u+t(v-u))+(y_{2}-y_{1})\frac{\partial f}{\partial y}(u+t(v-u)).$
However, on the domain of definition of $h$, we have $\frac{\partial f}{\partial x}>0$ and $\frac{\partial f}{\partial y}>0$. Therefore, $h^{\prime}(t)>0$ for all $t\in\left[0,1\right]$, which is impossible since $h$ is $\operatorname{\mathcal{C}^{1}}$ and by (14) and (15).
The proof of Theorem 4.2 follows the same approach as Theorem 4.1, with adjustments in the selection of vectors and the direction of the isoclines.
Proof of Theorem 4.2.
$\left(x^{*},y^{*}\right)$ is an attractive equilibrium point with the assumptions $\frac{\partial f}{\partial y}(x^{*},y^{*})>0$ and $\frac{\partial g}{\partial x}(x^{*},y^{*})>0$. According to Table 1 (see Appendix C), this case only arises when, in addition, $\frac{\partial f}{\partial x}(x^{*},y^{*})<0$ and $\frac{\partial g}{\partial y}(x^{*},y^{*})<0$. Since the functions $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$, there exists a neighborhood $R_{2}$ around the point $(x^{*},y^{*})$ such that $\frac{\partial f}{\partial y}>0$, $\frac{\partial g}{\partial x}>0$, $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$. Then both isoclines $\Gamma_{f}$ and $\Gamma_{g}$ are increasing. To correspond to an attractive equilibrium point (with index +1+1), the sign alternation around this point must match that of Figure 2(a). There are only two possible phase portrait configurations in the neighborhood see Figure 11.
(a) First possible alternation
(b) Second possible alternation
Figure 11: Illustration of Proof of Theorem 4.2. Two possible neighborhoods of the attractive equilibrium point $(x^{*},y^{*})$ that could lead to the formation of a repulsive limit cycle around it.
Let us consider the case of the Figure 11(a). The other case of the Figure 11(b) follows similarly with $g$. We know the sign of $f$ and $g$ on either side of $\Gamma_{f}$ and $\Gamma_{g}$. Let $u$ and $v$ be points as shown in Figure 11(a), with respective coordinates $\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix}$ and $\begin{pmatrix}x_{2}\\ y_{2}\end{pmatrix}$, where $x_{1}>x_{2}$ and $y_{1}<y_{2}$, and we define the function $h$ as in the previous proof, following (12), and obtain similar results as in (14) and (15). Since $h$ is $\operatorname{\mathcal{C}^{1}}$ and has the same derivative as in the previous proof, we arrive at the same contradiction.
We end with the proof of Theorem 4.3 and 4.4 which provide conditions under which stable limit cycles may be observed.
Proof of Theorem 4.3.
We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 5. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
AA lies along a trajectory, defined by $x=0$, $y>0$, and $(0,0)$ is an equilibrium point. Thus, no trajectory can escape across $A$ as $t\to+\infty$, because to do so, it would have to intersect another trajectory or $(0,0)$, and trajectories cannot intersect. $B$ consists of: two critical points, located at $(0,0)$ and $(x_{2},0)$, the trajectory defined by $0<x<x_{2}$, $y=0$, which moves away from $(0,0)$ towards the point $(x_{2},0)$, and finally the trajectory defined by $x_{2}<x$, $y=0$, which moves towards the point $(x_{2},0)$. As before, no trajectory can exit across $B$. Trajectories cross $C$ moving from right to left as $t$ increases, since the points along this arc are located above $\Gamma_{f}$. Given assumption 1.10 where $f<0$, this implies that $\dot{x}<0$ along the arc. The trajectory through the corner $e$ is horizontal and is moving to the left at that point, as $e$ lies on $\Gamma_{g}$, which means $\dot{y}=0$. Along $D$, trajectories move downward and to the left, since the points on this arc lie above both isoclines. Therefore, no trajectories can escape across the boundary of the rectangular region, and trajectories enter the region through the points along the top and right edges.
Our goal now is to demonstrate that the limit set $\omega(\gamma)$ of the trajectory $\gamma$ is a limit cycle. According to the fivefold way (see Coleman, (1983), page 265), this will follow if we can show that $\omega(\gamma)$ contains no critical points. The critical points of system (1) are located at the origin $(0,0)$, at the point $(x_{2},0)$ on the $x$-axis, and at the point $\left(x^{*},y^{*}\right)$. $\left(x^{*},y^{*}\right)$ cannot be in $\omega(\gamma)$ because $\left(x^{*},y^{*}\right)$ is repusilve by assumption 1.9. Moreover we cannot have that $\omega(\gamma)=\{(0,0)\}$. In fact, for the trajectory $\gamma$ to approach $(0,0)$, it must intersect the interior of the region where $f>0$ and $g<0$ (as shown in Figure 5), and stay in this region as $t\to+\infty$. However, if $\gamma$ is in the interior of this region, then $f>0$ and $\dot{x}>0$, which implies that we cannot have both $x(t)$, $y(t)$ remaining in the region for all large enough $t$ and $(x(t),y(t))\rightarrow(0,0)$ as $t\rightarrow+\infty$. A similar argument shows that $\omega(\gamma)\neq\{(x_{2},0)\}$.
Thus the last case is if $\omega(\gamma)$ contains any critical points at all, $\omega(\gamma)$ can be a cycle graph. $\left(x^{*},y^{*}\right)$ cannot belong to a cycle graph since $\left(x^{*},y^{*}\right)$ is not the $\omega$-limit set of any nonconstant trajectory. If $(x_{2},0)$ belongs to a cycle graph, then so must the orbit along $B$ with $\omega$-limit set $\{(x_{2},0)\}$ and $\alpha$-limit set $\{(0,0)\}$, since that is the only orbit in the rectangular region with $\omega$-limit set $\{(x_{2},0)\}$. But in a cycle graph, each limit set of each orbit is a critical point of the cycle graph. We must then have that $\{(0,0)\}$ in the cycle graph, then the unbounded orbit on $A$ also belongs to the cycle graph since it is the only orbit in the positive quadrant whose $\omega$-limit set is $\{(0,0)\}$. This is impossible since $\omega(\gamma)$ is bounded because $\gamma$ is bounded. Thus $\omega(\gamma)$ contains no critical points. By the fivefold way (see Coleman, (1983), page 265) we have that $\omega(\gamma)$ is a cycle. ∎
Proof of Theorem 4.4.
We will follow the same approach as in the previous proof. We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 7. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
AA consists of: two critical points, located at $(0,0)$ and $(0,y_{2})$, the trajectory defined by $0<y<y_{2}$, $x=0$, which moves away from $(0,0)$ towards the point $(0,y_{2})$, and finally the trajectory defined by $y_{2}<y$, $x=0$, which moves towards the point $(0,y_{2})$. Thus, no trajectory can escape across $A$ as $t\to+\infty$, because to do so, it would have to intersect another trajectory or $(0,0)$. $B$ lies along a trajectory, defined by $x>0$, $y=0$, and $(0,0)$ is an equilibrium point. As before, no trajectory can exit across $B$. Trajectories cross $C$ moving from right to left as $t$ increases, since the points along this arc are located below $\Gamma_{f}$. Given assumption 2.9 where $f<0$, this implies that $\dot{x}<0$ along the arc. The trajectory through the corner $e$ is vertical and is moving downwards at that point, as $e$ lies on $\Gamma_{f}$. Along $D$, trajectories move downward and to the right, since the points on this arc lie above both isoclines. Therefore, no trajectories can escape across the boundary of the rectangular region, and trajectories enter the region through the points along the top and right edges.
As in the previous proof, we will demonstrate that the limit set $\omega(\gamma)$ of the trajectory $\gamma$ is a limit cycle. It will follow if we can show that $\omega(\gamma)$ contains no critical points. The critical points of system (1) are located at the origin $(0,0)$, at the point $(0,y_{2})$ on the $y$-axis, and at the point $\left(x^{*},y^{*}\right)$. $\left(x^{*},y^{*}\right)$ cannot be in $\omega(\gamma)$ because $\left(x^{*},y^{*}\right)$ is repusilve by assumption 2.8. Moreover we cannot have that $\omega(\gamma)=\{(0,0)\}$. For the trajectory $\gamma$ to approach $(0,0)$, it must intersect the interior of the region where $f<0$ and $g>0$ (as shown in Figure 7), and stay in this region as $t\to+\infty$, which is impossible as before and a similar argument shows that $\omega(\gamma)\neq\{(0,y_{2})\}$.
Then the last case is if $\omega(\gamma)$ contains any critical points at all, $\omega(\gamma)$ can be a cycle graph. $\left(x^{*},y^{*}\right)$ cannot belong to a cycle graph as before. If $(0,y_{2})$ belongs to a cycle graph, then so must the orbit along $A$ with $\omega$-limit set $\{(0,y_{2})\}$ and $\alpha$-limit set $\{(0,0)\}$, since that is the only orbit in the rectangular region with $\omega$-limit set $\{(0,y_{2})\}$. But in a cycle graph, each limit set of each orbit is a critical point of the cycle graph. We must then have that $\{(0,0)\}$ in the cycle graph, then the unbounded orbit on $B$ also belongs to the cycle graph since it is the only orbit in the positive quadrant whose $\omega$-limit set is $\{(0,0)\}$. This is impossible since $\omega(\gamma)$ is bounded because $\gamma$ is bounded. Thus $\omega(\gamma)$ contains no critical points. By the fivefold way, we have that $\omega(\gamma)$ is a cycle. ∎
Acknowledgments
We sincerely thank Frédéric Faure for his valuable discussions and teachings. This work has benefited from the support of the MMB Chair (Mathematics for Life and Environmental Sciences) VEOLIA-Ecole Polytechnique-MNHN-F.X. The European Union (ERC Consolidator) grant OUTOFTHEBLUE (project number 101088089) obtained by VL supported this work. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Index Theory and Equilibrium Classification
Index theory is an interesting tool in the qualitative study of dynamical systems. It provides a way to classify equilibrium points and understand the possible global behavior of trajectories. The index of an equilibrium point is a topological invariant that remains unchanged under continuous deformations of the vector field.
A.1 Definition of the Index
Let $(x^{*},y^{*})$ be an equilibrium point of a planar dynamical system. The index of this equilibrium point is defined by analyzing the behavior of the vector field along a simple closed curve surrounding $(x^{*},y^{*})$. More formally, the index is given by the total winding number of the vector field as one traverses the closed curve counterclockwise.
Mathematically, a system given by:
$\left\{\begin{array}[]{ll}\dot{x}&=F\left(x,y\right)\\ \dot{y}&=G\left(x,y\right)\end{array}\right$. (16)
defines a vector field for which the slope:
$\frac{dy}{dx}=\frac{G(x,y)}{F(x,y)}$ (17)
forms an angle with the $x$-axis given by:
$\varphi(x,y)=\arctan\left(\frac{G(x,y)}{F(x,y)}\right).$ (18)
Given this angle and a simple closed curve $\gamma$ in $\mathbb{R}^{2}$, the index of $\gamma$ , denoted $\text{Ind}(\gamma)$, is defined as:
$\text{Ind}(\gamma)=\frac{1}{2\pi}\oint_{\gamma}d\varphi(x,y).$ (19)
A.2 Indices of Different types of Equilibrium Points
The index of an equilibrium point depends on its stability and the local structure of trajectories.
  • Saddle Point (Index 1-1): A saddle point is an equilibrium where trajectories approach along one direction and diverge along another. The vector field rotates by $\pi$ as one completes a loop around the equilibrium, leading to an index of 1-1.
  • Attractive Point (Index +1+1): An attractor is an equilibrium where all nearby trajectories converge. The vector field makes a full $2\pi$ rotation around the equilibrium, giving an index of +1+1.
  • Repulsive Point (Index +1+1): A repeller is an equilibrium where all nearby trajectories move away. As with attractors, the vector field completes a full rotation around the point, resulting in an index of +1+1.
A.3 Implications of Index Theory
A key consequence of index theory is the Poincaré-Hopf theorem, which states that the sum of the indices of all equilibrium points within a bounded region must equal the Euler characteristic of the domain. This imposes constraints on the possible configurations of equilibria and their stability.
For example, in a dynamical system where the entire phase plane is enclosed within a region, the total index must sum to +1+1. This means that a system with a single repulsive equilibrium point cannot have only saddle points without additional attractors or limit cycles to balance the total index.
Appendix B Trace-Determinant Analysis: Classical study of the point $(0,0)$
The system (1) always has two isoclines ($y=0$ and $x=0$), making $(0,0)$ an equilibrium point. To determine its nature, we examine the Jacobian matrix at $(0,0)$, assuming $f(0,0)\neq 0$ and $g(0,0)\neq 0$:
$J(0,0)=\begin{pmatrix}f(0,0)&0\\ 0&g(0,0)\end{pmatrix}$.
with trace $\text{Tr}(J)=f(0,0)+g(0,0)$ and determinant $\text{det}(J)=f(0,0)g(0,0)$. Three cases arise:
  • (i)
    Repulsive node ($f(0,0)>0$, $g(0,0)>0$): Both species grow at low densities, leading to population increase. This suggests facultative mutualism or independent persistence. In phase portrait, trajectories move away from extinction.
  • (ii)
    Attractive node ($f(0,0)<0$, $g(0,0)<0$): Both species decline at low densities, leading to extinction. It can also indicate obligate mutualism, where each species is highly dependent on the other, and without sufficient partner density, both face extinction. This may indicate obligate mutualism or Allee effects. In phase portrait, trajectories converge to extinction.
  • (iii)
    Saddle point ($f(0,0)g(0,0)<0$): One species can persist alone, while the other requires interaction to survive. This may indicate an asymmetric dependency (facultative-obligate mutualism where one species depends on the other for survival). Some trajectories lead to extinction, while others promote coexistence or growth
These cases determine whether extinction is stable or unstable and how interactions shape population dynamics at low densities.
Appendix C Isocline behaviors and Point classifications
[…]
$\frac{\partial g}{\partial x}>0$ $\frac{\partial g}{\partial x}>0$ $\frac{\partial g}{\partial x}<0$ $\frac{\partial g}{\partial x}<0$
$\frac{\partial g}{\partial y}>0$ $\frac{\partial g}{\partial y}<0$ $\frac{\partial g}{\partial y}>0$ $\frac{\partial g}{\partial y}<0$
$\frac{\partial f}{\partial x}>0$ Repulsive or Saddle Point Saddle Point Repulsive Indeterminate
$x$-isocline $\searrow$ $x$-isocline $\searrow$ $x$-isocline $\searrow$ $x$-isocline $\searrow$
$\frac{\partial f}{\partial y}>0$ $y$-isocline $\searrow$ $y$-isocline $\nearrow$ $y$-isocline $\nearrow$ $y$-isocline $\searrow$
Strict Mutualism Strict Mutualism
$\frac{\partial f}{\partial x}>0$ Repulsive Indeterminate Repulsive or Saddle Point Saddle Point
$x$-isocline $\nearrow$ $x$-isocline $\nearrow$ $x$-isocline $\nearrow$ $x$-isocline $\nearrow$
$\frac{\partial f}{\partial y}<0$ $y$-isocline $\searrow$ $y$-isocline $\nearrow$ $y$-isocline $\nearrow$ $y$-isocline $\searrow$
$\frac{\partial f}{\partial x}<0$ Saddle Point Attractive or Saddle Point Indeterminate Attractive
$x$-isocline $\nearrow$ $x$-isocline $\nearrow$ $x$-isocline $\nearrow$ $x$-isocline $\nearrow$
$\frac{\partial f}{\partial y}>0$ $y$-isocline $\searrow$ $y$-isocline $\nearrow$ $y$-isocline $\nearrow$ $y$-isocline $\searrow$
Strict Mutualism Strict Mutualism
$\frac{\partial f}{\partial x}<0$ Indeterminate Attractive Saddle Point Attractive or Saddle Point
$x$-isocline $\searrow$ $x$-isocline $\searrow$ $x$-isocline $\searrow$ $x$-isocline $\searrow$
$\frac{\partial f}{\partial y}<0$ $y$-isocline $\searrow$ $y$-isocline $\nearrow$ $y$-isocline $\nearrow$ $y$-isocline $\searrow$
Table 1: Summary of isocline behaviors and point classifications.

### v2

@@
 1 Introduction
-In natural communities, interactions between individuals from different species are widespread, ranging from antagonistic relationships, such as competition or parasitism, to cooperative ones, such as mutualism (Bronstein,, 2015; Janzen,, 1985).
-Mutualistic interactions include processes such as seed dispersal (Janzen,, 1985), pollination (Klein et al.,, 2007), resource exchange (van der Heijden et al.,, 2008), and protection (Bronstein,, 2015), where individuals from both species may benefit from the relationship.
-For example, mycorrhizal fungi provide plants with essential nutrients, while plants provide carbohydrates to the interacting fungi (Bonfante and Genre,, 2010).
-Similarly, in pollination, plants produce nectar feeding some insects, which in turn transport pollen and facilitate plant reproduction (Sargent and Ackerly,, 2008).
-Mutualistic interactions can be either facultative or obligatory.
-If facultative, the dynamics of each species can benefit from each other, but each species can be maintained in the absence of the other (e.g.
-generalist pollination, where plants and pollinators can often find alternative partners (Sargent and Ackerly,, 2008)).
-Otherwise, if mutualistic interactions are obligatory, the persistence of each species entirely depends on the mutualistic relationship and hence on the presence of the interacting species (e.g.
-the fig-wasp relationship, Bronstein, 2001).
-The nature and strength of the interactions can also depend on population densities.
-Variations in species densities can shift the cost-benefit dynamics of the relationship, consequently shifting interactions along a continuum from mutualism to parasitism.
+In natural communities, ecological interactions between individuals from different species are widespread and profoundly shape the dynamics of populations.
+These interactions can be either antagonistic relationships, such as competition or parasitism, or cooperative ones, such as mutualism (Bronstein, 2015; Janzen, 1985), therefore leading to different demographic outcomes in the populations involved.
+Mutualistic interactions, where individuals from interacting species may benefit from the relationship, include processes such as seed dispersal (Janzen, 1985), pollination (Klein et al., 2007), resource exchange (van der Heijden et al., 2008), and protection (Bronstein, 2015).
+For example, mycorrhizal fungi provide plants with essential nutrients, while plants provide carbohydrates to the interacting fungi (Bonfante and Genre, 2010).
+Similarly, in pollination, plants produce nectar feeding some insects, which in turn transport pollen and facilitate plant reproduction (Sargent and Ackerly, 2008).
+Mutualistic relationships can be either facultative or obligatory, hence triggering different consequences on the co-existence and dynamics of interacting populations.
+If mutualistic interactions are obligatory, the persistence of each species entirely depends on the mutualistic relationship and hence on the presence of the interacting species (e.g.
+the fig-wasp specialized relationship (Bronstein, 2001)).
+Otherwise, in facultative interactions, the dynamics of each species can benefit from each other, but each species can be maintained in the absence of the other (e.g.
+generalist pollination, where plants and pollinators can often find alternative partners (Sargent and Ackerly, 2008)).
+However, the benefits brought to the individuals from the interacting species are not necessarily symmetrical nor constant.
+While the mutualistic interactions are classically assumed to bring only benefits to the individuals involved, they can also bring costs, especially in certain conditions of respective densities.
+Density-dependent benefits have indeed been documented in several communities.
+Respective densities of interacting species can shift the cost-benefit dynamics of the relationship, consequently shifting interactions along a continuum from mutualism to parasitism.
 For example, in ant-plant mutualism, plants provide ants with food in exchange for protection.
-However, at low ant densities, plants may not receive sufficient protection, while at high densities, the cost of supporting the ants may outweigh the protective benefits (Bronstein,, 1994; Holland et al.,, 2002; Janzen,, 1985).
-Similarly, in pollination systems, the investment of plants in nectar can become inefficient if pollinator densities are too low to ensure effective pollination, or too high, leading to increased competition among pollinators and diminishing returns for the plants (Kawai and Tokeshi,, 2007; Morris et al.,, 2010; Holland et al.,, 2002).
-Mathematical models of ecological interactions have mostly focused on antagonistic relationships such as predator-prey and competitive dynamics.
-While some models addressed mutualism, they often assumed strictly positive interactions, leading to unrealistic outcomes like unbounded population growth or infinite densities (Brauer and Castillo-Chavez,, 2012; Hale and Valdovinos,, 2021; May,, 1972).
-Indeed, because of the complexity and context-dependence of mutualistic interactions, modeling them as a direct extension of Lotka–Volterra systems overlooks key mechanisms, such as density-dependent saturation, that are essential to preventing unbounded growth.
-Moreover, the factors shaping the persistence of mutualistic interactions and the transitions between parasitic and mutualistic interactions driven by populations densities variations (i.e., on ecological time scales) are much less documented than in predator–prey systems (Holland et al.,, 2002; Revilla and Encinas-Viso,, 2015).
-Density dependence and saturation effects to prevent unbounded growth or systematic extinction were considered in Brauer and Castillo-Chavez, (2012), bounded growth terms were added in Holland et al., (2002) to model that mutualistic benefits naturally diminish as population sizes increase, and finally Ruiz-Herrera, (2025) analyzed mutualistic systems with a density-dependent framework, emphasizing the role of benefit–cost ratios in shaping long-term interaction outcomes.
+However, at low ant densities, plants may not receive sufficient protection, while at high densities, the cost of supporting the ants may outweigh the protective benefits (Bronstein, 1994; Holland et al., 2002; Janzen, 1985).
+Similarly, in pollination systems, the investment of plants in nectar can become inefficient if pollinator densities are too low to ensure effective pollination, or too high, leading to increased competition among pollinators and diminishing returns for the plants (Kawai and Tokeshi, 2007; Morris et al., 2010; Holland et al., 2002).
+There is thus a wide diversity of mutualistic interactions in the wild, and the dynamics of the different species involved can differ depending on the ecological context.
+Their consequences on population dynamics have been previously described mostly by mathematical models focusing on specific ecological situations, therefore limiting general predictions on the effects of mutualistic interactions.
+Many classical models addressing mutualism assumed strictly positive interactions, leading to unrealistic outcomes like unbounded population growth or infinite densities (Brauer and Castillo-Chavez, 2012; Hale and Valdovinos, 2021; May, 1972).
+Indeed, because of the complexity and context-dependence of mutualistic interactions, modeling them as a direct extension of Lotka–Volterra systems overlooks key mechanisms, such as density-dependent saturation, that prevent unbounded growth in actual populations.
+Moreover, the factors shaping the persistence of mutualistic interactions and the transitions between parasitic and mutualistic interactions driven by populations densities variations (i.e.
+on ecological time scales) are poorly documented in mutualistic interactions, while they are well-known in predator–prey systems modeled using Lokta-Volterra equations (Holland et al., 2002; Revilla and Encinas-Viso, 2015).
+Brauer and Castillo-Chavez (2012) and Holland et al.
+(2002) considered mutualistic models with saturation effects preventing either unbounded growth or systematic extinction.
+More recently Ruiz-Herrera (2025) analyzed mutualistic systems with a density-dependent framework, emphasizing the role of benefit–cost ratios in shaping long-term interaction outcomes.
 Asymmetries in obligate vs.
-facultative mutualistic interactions among species have also been shown to have a significant influence on system stability and resilience (Bronstein,, 1994; Holland and DeAngelis,, 2010; Neuhauser and Fargione,, 2004).
-Most of these previous models studying the dynamics of mutualistic interactions rely on ordinary differential equations to describe the conditions for coexistence and stability (see Hale and Valdovinos, (2021) for a review).
-Other approaches such as network analysis have also been used to explore interspecific relationships within ecological systems (Bascompte et al.,, 2003) and partial differential equations to model the spatial dynamics of mutualistic interactions (Aliyu and Mohd,, 2021).
-Examples of mutualistic interactions with density-dependent effect have been documented in natural populations (Bronstein,, 2015).
-Yet, no general theory has been developed to study the coexistence of mutualistic species while accounting for this important process.
-As pointed out by Hale and Valdovinos, (2021), it is striking that all mutualistic interactions models studied so far show similar outcomes and behaviors, suggesting that some general properties can be found, independently of the peculiarities of each model.
+facultative mutualistic interactions among species have also been shown to significantly affect system stability and resilience in a series of models (Bronstein, 1994; Holland and DeAngelis, 2010; Neuhauser and Fargione, 2004).
+Most of these previous models studying the dynamics of mutualistic interactions relied on ordinary differential equations to describe the conditions for coexistence and stability (see Hale and Valdovinos (2021) for a review).
+Other approaches such as network analysis have also been used to explore interspecific relationships within ecological systems (Bascompte et al., 2003) and partial differential equations to model the spatial dynamics of mutualistic interactions (Aliyu and Mohd, 2021).
+Despite this diversity of modeling approaches, a general theoretical framework investigating the coexistence and dynamics of mutualistic species while accounting for any type of density-dependent effects is still lacking.
+At the same time, as emphasized by Hale and Valdovinos (2021), it is striking that all mutualistic interaction models studied so far show similar outcomes and behaviors, suggesting that some general properties can be found, independently of the peculiarities of each model.
 For example, there must be some general conditions for the stable coexistence of the two species when the interaction is density-dependent.
@@
 While these models provide valuable insights, they rarely consider the possibility of more complex dynamics, such as limit cycles or oscillatory behaviors, therefore preventing general predictions.
-By proposing and analyzing a general mathematical model of two-species mutualism, we aim at tackling these limitations, focusing in particular on two main questions:
-  1. 1.
-    What types of dynamical behaviors, including equilibrium configurations, can be derived from models of mutualistic interactions, and which types are not covered by these models?
-  2. 2.
-    To what extent does incorporating transitions between mutualism and parasitism allow us to observe previously neglected dynamics?
-To address these questions, Section 2 first provides the general definitions of density-dependent effects acting on the dynamics of species involved in ecological interactions, which will be used throughout the paper.
-We thus introduce a system of differential equations for two interacting populations, incorporating density-dependent interactions, without focusing on specific types of interactions within a particular ecological context and on limited range of parameters.
+By building and analyzing a mathematical model of two-species mutualism with general equations allowing to cover any kind of density-dependent processes, we aim to overcome these limitations and to clarify the range of dynamical behaviors that can emerge in mutualistic systems, including equilibrium structures.
+To this end, we adopt a framework that is general in two complementary senses.
+First, the model proposed here is mathematically general: instead of prescribing specific functional responses or mechanistic assumptions, we rely on qualitative and structural constraints on the growth functions and the signs of their partial derivatives.
+This enables the model to encompass a wide class of two-species interaction systems, including but not limited to most previously published mutualistic models (see Hale and Valdovinos (2021) and Tables in Section 2.1).
+Second, the model described here is biologically general: because interactions are represented through density-dependent signs of the cross-partial derivatives, the framework is able to incorporate a broad spectrum of ecological processes such as mutualistic benefits, saturation, costs, and shifts to parasitic effects at high densities.
+Thus, it provides a unified tool for ecological relationships that do not necessarily remain strictly mutualistic across all possible population densities.
+This paper is structured as follows: Section 2 first provides the general definitions of density-dependent effects acting on the dynamics of species involved in ecological interactions, which will be used throughout the paper.
+We introduce a system of differential equations for two interacting populations, incorporating density-dependent interactions, without focusing on specific types of interactions within a particular ecological context and on limited range of parameters.
 In Section 3, we formulate general and robust assumptions derived from the analysis of deterministic mutualistic models (Theorem 3.1).
@@
 Because we model population sizes, the analysis is confined to the non-negative quadrant of $\mathbb{R}^{2}$.
-This form is commonly used to describe two-species dynamics without accounting for immigration, as in the models of Hale and Valdovinos, (2021), May, (1972), and Brauer and Castillo-Chavez, (2012).
+This form is commonly used to describe two-species dynamics without accounting for immigration (see Thompson et al.
+(2006)), as in the models studied in May (1972), Brauer and Castillo-Chavez (2012), and reviewed in Hale and Valdovinos (2021).
+The sections 2.1.1 to 2.1.3 present examples of mutualism models studied in previous works, which are of the form (1).
+For the sake of readability, we have divided them into three groups according to their underlying ecological mechanisms.
+They are presented with their corresponding $f$ and $g$ functions from Eq.
+(1).
+2.1.1 Linear–Benefit Mutualism Models
+Table 1 compiles mutualism models where benefit increases linearly with partner density and there is no saturation effect.
+Model Per-capita growth functions Interpretation Reference
+Linear benefit with logistic limitation $\begin{aligned} f(x,y)&=r_{1}\left(\dfrac{K_{1}+\alpha_{12}y-x}{K_{1}}\right)\\ g(x,y)&=r_{2}\left(\dfrac{K_{2}+\alpha_{21}x-y}{K_{2}}\right)\end{aligned}$ Linear increase of partner benefit; modification of intrinsic growth.
+Gause and Witt (1935)
+Symbiont–host linear model $\begin{aligned} f(x,y)&=r_{1}\frac{K_{1}+\alpha_{12}y-x}{K_{1}+\alpha_{12}y}\\ g(x,y)&=r_{2}\frac{K_{2}+\alpha_{21}x-y}{K_{2}}\end{aligned}$ Asymmetric symbiont–host structure; linear modification of carrying capacity; obligate for $K_{1}=0$.
+Whittaker (1970)
+Lotka–Volterra mutualism $\begin{aligned} f(x,y)&=r_{1}+\alpha_{12}y-\alpha_{11}x\\ g(x,y)&=r_{2}+\alpha_{21}x-\alpha_{22}y\end{aligned}$ Classical linear mutualistic enhancement of per-capita growth; obligate when $r_{1}/\alpha_{11}\leq 0$.
+Vandermeer and Boucher (1978)
+Multiplicative linear benefit $f(x,y)$ $=r_{1}\left(1-\dfrac{x}{K_{1}}\right)\left(1+\dfrac{\alpha_{12}y}{K_{1}}\right)$ $g(x,y)$ $=r_{2}\left(1-\dfrac{y}{K_{2}}\right)\left(1+\dfrac{\alpha_{21}x}{K_{2}}\right)$ Partner increases effective carrying capacity multiplicatively; facultative.
+Addicott (1981)
+Table 1: Linear–benefit mutualism models expressed through per-capita growth functions $f$ et $g$.
+2.1.2 Saturating Mutualism Models (Interspecific Saturation)
+Table 2 compiles mutualism models where the benefits gained by a specie is a saturating function of the density of the other species.
+Model Per-capita growth functions Interpretation Reference
+Birth limitation reduced by partner $\begin{aligned} f(x,y)&=r_{1}-\frac{b_{1}x}{1+m_{1}y}-d_{1}x\\ g(x,y)&=r_{2}-\frac{b_{2}y}{1+m_{2}x}-d_{2}y\end{aligned}$ Partner reduces density-dependent birth limitation; saturating effect with partner density; facultative.
+Wolin and Lawlor (1984)
+Saturating forager mutualism $\begin{aligned} f(x,y)&=r_{1}-c_{1}x+\frac{b_{12}a_{12}y}{1+a_{12}h_{12}y}\\ g(x,y)&=r_{2}-c_{2}y+\frac{b_{21}a_{21}x}{1+a_{21}h_{21}x}\end{aligned}$ Handling time limits mutualistic gains; stabilizes coexistence; fits pollination/foraging systems.
+Wright (1989)
+Exponential saturation of benefits $f(x,y)$ $=r_{10}+(r_{11}-r_{10})$ $\times\left(1-e^{-k_{1}y}\right)-a_{1}x$ $g(x,y)$ $=r_{20}+(r_{21}-r_{20})$ $\times\left(1-e^{-k_{2}x}\right)-a_{2}y$ Benefits saturate exponentially with partner density.
+Graves et al.
+(2006)
+Plant–pollinator with shared saturation $\begin{aligned} f(x,y)&=\frac{\eta\alpha y}{1+\alpha x+\alpha\beta y}-b-cx\\ g(x,y)&=\frac{\mu\alpha x}{1+\alpha x+\alpha\beta y}-d\end{aligned}$ Shared saturation of benefits; obligate mutualism.
+Fishman and Hadany (2010)
+Table 2: Mutualism models with interspecific saturation of benefits expressed through per-capita growth functions $f$.
+2.1.3 Non-Monotonic Mutualism Models and Consumer-Resource Approach
+Table 3 compiles mutualism models where the benefits gained by a specie is a non-monotonic function of the density of the other species.
+Model Per-capita growth functions Interpretation Reference
+Quadratic interaction $\begin{aligned} f(x,y)&=r_{1}\left(c_{1}-x-a_{1}(y-b_{1})^{2}\right)\\ g(x,y)&=r_{2}\left(c_{2}-y-a_{2}(x-b_{2})^{2}\right)\end{aligned}$ Benefit maximized at intermediate partner density; allows transitions between mutualism and competition.
+Zhang (2003)
+Plant–mycorrhizae with cost $\begin{aligned} f(x,y)&=r_{1}\frac{K_{1}+\alpha_{12}y-x}{K_{1}+\alpha_{12}y}-ay\\ g(x,y)&=r_{2}\frac{K_{2}+\alpha_{21}x-y}{K_{2}}\end{aligned}$ Interspecific asymmetric cost; facultative mutualism.
+Neuhauser and Fargione (2004)
+Protection mutualism with saturating mortality reduction $f(x,y)$ $=\rho_{1}b_{1}\left(1-\frac{x}{S_{1}}\right)$ $-\left(d_{1\min}+\frac{d_{1\max}-d_{1\min}}{1+c_{1}y}\right)$ $g(x,y)$ $=\rho_{2}b_{2}\left(1-\frac{y}{S_{2}+x}\right)$ $-\left(d_{2\min}+\frac{d_{2\max}-d_{2\min}}{1+c_{2}x}\right)$ Protection reduces mortality at low partner density; net effects may shift with density.
+Thompson et al.
+(2006)
+Consumer–resource mutualism $\begin{aligned} f(x,y)&=r_{1}+c_{1}\frac{a_{12}y}{h_{2}+y}\\ &-q_{1}\frac{\beta_{12}y}{e_{1}+x}-s_{1}x\\ g(x,y)&=r_{2}+c_{2}\frac{a_{21}x}{h_{1}+x}\\ &-q_{2}\frac{\beta_{21}x}{e_{2}+y}-s_{2}y\end{aligned}$ Bidirectional consumer–resource structure; obligate when $r_{1}=0$.
+Holland and DeAngelis (2010)
+Ant–fungal garden mutualism $\begin{aligned} f(x,y)&=r_{f}\frac{ay^{2}}{b+ay^{2}}-r_{c}y-d_{1}x\\ g(x,y)&=r_{a}x-d_{2}y\end{aligned}$ Consumer–resource mutualism with nonlinear resource production.
+Kang et al.
+(2011)
+Plant–mycorrhizae with variable costs $f(x,y)$ $=r_{p}+q_{hp}\frac{\alpha y}{d+x}$ $-q_{cp}\beta y-\mu_{p}x$ $g(x,y)$ $=q_{cm}\beta x-q_{hm}\frac{\alpha x}{d+x}-\mu_{m}y$ Explicit separation of benefit and cost terms; obligate mutualism.
+Martignoni et al.
+(2020)
+Plant–pollinator consumer–resource model $f(x,y)$ $=b_{P}\left(f+\phi\frac{axy}{1+ahx+axy}\right)$ $-s_{P}x-d_{P}$ $g(x,y)$ $=b_{A}+\epsilon\frac{ax}{1+ahx}-s_{A}y-d_{A}$ Mechanistic pollination model.
+Hale and Valdovinos (2021)
+Table 3: Non-monotonic mutualism and consumer-resource models with peak benefit at intermediate partner density expressed through per-capita growth functions $f$.
 2.2 Key Assumptions and Definition
-To analyze and interpret the behavior of the system, we follow an approach similar to that used in Brauer and Castillo-Chavez, (2012), relying on the partial derivatives of the growth functions.
+We have so far not explicitly defined what makes a mutualistic model different than a model considering predator–prey or competitive interactions.
+The system of differential equations (1) does not, by itself, specify the nature of the interactions, as it also includes classical models such as the Lotka–Volterra predator–prey system.
+We introduce hereafter definitions of mutualism, parasitism, competition, each imposing specific constraints on the functional forms of $f$ and $g$.
+To analyze and interpret the behavior of the general system (1) in a mutualistic context, we follow an approach similar to Brauer and Castillo-Chavez (2012), relying on the partial derivatives of the growth functions.
 These derivatives capture how the growth rate of each species responds to changes in the density of both species.
@@
 We focus on specific regions of the phase plane defined by the population densities $x$ and $y$.
-In each region, the sign of the partial derivatives of the interaction terms determines the local effect of one species on the growth rate of the other.
+In each region, the signs of the partial derivatives of the interaction terms determine the local effect of one species on the growth rate of the other.
 As a result, an interaction may be positive in one part of the phase portrait and negative in another, depending on the local population densities.
@@
   • Intraspecific Cooperation: $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ describe positive feedback within each species, representing cooperative interactions, where individuals of the same population support each other’s growth.
+This mechanism is also called ’Allee effect’ when the positive feedback occurs at low density.
   • Dynamic Transitions: The signs of $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ are not fixed and may change with species densities, reflecting a continuum of interactions ranging from mutualism to parasitism.
 The phase portrait may contain regions where the signs of $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ differ, reflecting changes in interaction type.
-In contrast to many pre-existing models where these partial derivatives always remain within the same strict sign or even remain constant regardless of species densities (Hale and Valdovinos,, 2021), we allow the sign of these partial derivatives to change.
+In contrast to many pre-existing models where these partial derivatives always remain within the same strict sign or even remain constant regardless of species densities (Hale and Valdovinos, 2021), we allow the sign of these partial derivatives to change.
 Such assumption allows for transitions from mutualistic (positive value) to parasitic interactions (negative value) between species, depending on their respective densities.
@@
 In the first case, a strictly mutualistic region appears, although the condition does not hold for all densities $x$ and $y$.
-In the second case, the system remains mutualistic overall, but no strictly mutualistic region exists; that is, both populations benefit from the interaction, though not simultaneously.
+In the second case, the system remains mutualistic overall, but no strictly mutualistic region exists; that is, both populations may benefit from the interaction, though not simultaneously.
 (a) Local region of strict mutualism: Overlapping positive partial derivatives.
@@
 $\frac{\partial f}{\partial y}>0$).
-Outside these regions, the interaction is locally parasitic.
+Outside these regions, the interaction is parasitic for both species.
 3 General dynamics of two species with any types of interactions
 In this work, we study the behavior of system (1), including stability, oscillations, and other dynamics, with a particular focus on how the specific forms of $f$ and $g$ influence species coexistence.
-In Section 3.1 we consider a general framework in which $f$ and $g$ are assumed to be continuous only, and we apply index theory (see Perko, (2013) and Appendix A for details) to prove Theorem 3.1.
-Then, using our general Theorem 3.1, we study the different assumptions described in Section 2.2 adding the assumption that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$ to investigate more precisely how the nature of the interactions considered can affect the outcomes of the system.
+In Sections 3.1, 3.2 and 3.3 we consider a general framework in which $f$ and $g$ are assumed to be continuous only, and we apply index theory (see Perko (2013) and Appendix A for details) to prove Theorem 3.1.
+Then, in Section 3.4, using our general Theorem 3.1, we more precisely investigate how the nature of the interactions can affect the outcomes of the system under the additional assumption that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$.
 We mainly focus on simple equilibrium points and use methods based on the Jacobian matrix to analyze their stability and nature.
-For example, regardless of the specific nature of the interaction represented by system (1), certain information can already be deduced about the equilibrium point $\left(0,0\right)$ by applying these methods (see Appendix B and Brauer and Castillo-Chavez, (2012)).
-3.1 Assumptions and Equilibrium Index Alternation
+For example, regardless of the specific nature of the interaction represented by system (1), certain information can already be deduced about the equilibrium point $\left(0,0\right)$ by applying these methods (see Appendix B and Brauer and Castillo-Chavez (2012)).
+3.1 Assumptions
 We introduce several assumptions about the functions $f$ and $g$ that are commonly derived from biological models of population dynamics.
-Most of these assumptions, inspired by models reviewed in Hale and Valdovinos, (2021), are robust and remain valid even when $f$ and $g$ are subject to minor local perturbations.
+Most of these assumptions, inspired by models reviewed in Hale and Valdovinos (2021), are robust and remain valid even when $f$ and $g$ are subject to minor local perturbations.
 We recall that the zeros of $f$ and $g$ define the isoclines of the system, and their intersections determine the equilibrium points.
@@
 These assumptions directly influence the properties of the vector field.
+3.2 Equilibrium Index Alternation
 One characteristic of equilibrium points is their index, which describes the local behavior of the flow in their neighborhood.
-The index helps determine whether an equilibrium point acts as an attractive, a repulsive, or a saddle point, and reveals how trajectories interact with the point (see Perko, (2013) and Appendix A for more details).
+The index helps determine whether an equilibrium point acts as an attractive, a repulsive, or a saddle point, and reveals how trajectories interact with the point (see Perko (2013) and Appendix A for more details).
 In particular, for systems satisfying H1 to H4 on $f$ and $g$, the indices of equilibrium points in the positive quadrant exhibit a regular alternation along the isoclines.
@@
 Remark 3.3.
-We recall that the model cover the wide parasitic–mutualistic continuum.
+We recall that the model covers the wide parasitic–mutualistic continuum.
 The interactions may therefore represent competition, predation, or mutualism, as long as the functions $f$ and $g$ satisfy the conditions stated in assumptions H1H4.
@@
 The figure illustrates a mutualistic system with multiple types of equilibrium points.
+3.3 Generalisation to Axes
+Theorem 3.1 describes the alternation of indices for equilibrium points located strictly inside the positive quadrant, i.e.
+at intersections of $\Gamma_{f}$ and $\Gamma_{g}$ with strictly positive coordinates.
+However, equilibrium points may also lie on the axes.
+These additional equilibrium points may arise from the intersections of $\Gamma_{f}$ with the axis $x=0$, or of $\Gamma_{g}$ with the axis $y=0$.
+For population models, such equilibria correspond to situations in which one species persists at equilibrium while the other is extinct.
+For isoclines $\Gamma_{f}$ and $\Gamma_{g}$ satisfying hypotheses H1 to H4, a finite number of geometric configurations may occur along the coordinate axes.
+Since each isocline is continuous and consists of a single connected component, it can intersect the axes at most twice.
+Taking into account the relative ordering of these intersections along the axes, a total of $71$ distinct configurations are possible.
+Among them, $18$ configurations generate at least one equilibrium point located on an axis and are therefore not covered by Theorem 3.1.
+Under the same geometric assumptions on the isoclines, this alternation property can be extended to include such boundary equilibria.
+Proposition 3.5.
+Under the assumptions of Theorem 3.1, the alternation of equilibrium point indices along the isoclines extends to include equilibrium points located on the coordinate axes, with the exception of the origin $(0,0)$.
+This result shows that the index alternation is a global geometric property of the vector field induced by (1), and is not restricted to strictly positive coexistence equilibria.
+Equilibrium points located on the axes naturally fit into the same alternation pattern, while the origin $(0,0)$ plays a singular role and is therefore excluded from this statement.
 3.4 Intraspecific Competition and Cooperation Assumption
@@
 These results are summarized in Table 4.
-In many models presented in Hale and Valdovinos, (2021), intraspecific competition is a key mechanism in regulating population dynamics within each species, preventing the uncontrolled population growth which may be criticized in some models of mutualistic interactions.
+In many models presented in Hale and Valdovinos (2021), intraspecific competition is a key mechanism in regulating population dynamics within each species, preventing the uncontrolled population growth which may be criticized in some models of mutualistic interactions.
 With this additional assumption, all previous equilibrium points with an index of +1+1 become attractive points.
@@
 For a model satisfying the conditions of Corollary 3.6, if there are at least two equilibrium points in the positive quadrant, then an attractive point exists.
-This implies the existence of conditions under which coexistence of the two species is possible.
+This implies the existence of conditions under which the coexistence of the two species is possible.
+Thus, negative assumptions on the sign of the partial derivatives $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$ (which biologically correspond to intraspecific competition, always negative regardless of species densities) prevent the emergence of any coexistence solutions other than attractive equilibrium points or saddle points.
+As in Proposition 3.5, this corollary can be extended to the coordinate axes, where the same monotonicity assumptions guarantee uniqueness and attractivity of equilibria.
+Corollary 3.8.
+Under the assumptions of Corollary 3.6, if an equilibrium point exists on a coordinate axis in $\mathbb{R}_{+}\times\mathbb{R}_{+}$, then it is unique on that axis.
+Moreover, if this equilibrium has index +1+1, then it is an attractive node.
 Under the opposite constant sign assumptions, i.e.
-intraspecific cooperation for both species as described in 2.2, the system in the positive quadrant alternates between repulsive nodes and saddle points along the isoclines.
+intraspecific cooperation for both species as described in Section 2.2, the system in the positive quadrant alternates between repulsive nodes and saddle points along the isoclines.
 Corollary 3.9.
@@
 Under these assumptions, this reflects a fragile coexistence scenario where small perturbations can push the system away from equilibrium, leading to instability in the dynamics of the two species.
-Thus, negative assumptions on the sign of the partial derivatives $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$ (which biologically correspond to intraspecific competition, always negative regardless of species densities) prevent the emergence of any coexistence solutions other than attractive equilibrium points or saddle points.
-3.3 Illustrative Examples of Theorem 3.1
-In this section, we illustrate Theorem 3.1 using four models of mutualistic interactions reviewed in Hale and Valdovinos, (2021), each represented on Figure 4 (in the order presented), each emphasizing a different characteristic of mutualism.
-These four models use different mathematical representations to emphasize certain aspects of mutualism over others, and describe facultative (Neuhauser and Fargione,, 2004) or obligate (Graves et al.,, 2006), density-dependent, or saturating mutualistic interactions.
+This corollary can also be extended to the coordinate axes: the same monotonicity properties imply that any equilibrium lying on an axis is unique there, and if it has index (+1), it is a repulsive equilibrium.
+Corollary 3.10.
+Under the assumptions of Corollary 3.9, if an equilibrium point exists on a coordinate axis in $\mathbb{R}_{+}\times\mathbb{R}_{+}$, then it is unique on that axis.
+Moreover, if this equilibrium has index +1+1, then it is a repulsive node.
+3.5 Illustration of Theorem 3.1 with Mutualistic Models from the Literature
+In this section, we illustrate Theorem 3.1 using four models of mutualistic interactions reviewed in Hale and Valdovinos (2021) (see Table 2 and Table 3), each represented on Figure 4 (in the order presented), each emphasizing a different characteristic of mutualism.
 Although their mathematical formulation reflects different perspectives on mutualistic interactions, as discussed in Section 2.2, we show that they satisfy the assumptions of system (1), highlighting the generality of the framework provided in our study.
   • (a)
-    In Zhang, (2003) the author models the dynamics of two mutualistic species using the same mathematical formulation twice, reflecting an interaction between species at the same trophic level.
+    In Zhang (2003) the author models the dynamics of two mutualistic species using the same mathematical formulation twice, reflecting an interaction between species at the same trophic level.
 In this model, the growth rate of each species is influenced by a quadratic term that depends on the density of the other species.
@@
 This formulation captures the effect of saturation and illustrates the continuum between mutualism and parasitism.
-The system is given by:
-    $\begin{cases}\dot{x}=r_{x}x\left(c_{x}-x-a_{x}(y-b_{x})^{2}\right),\\ \dot{y}=r_{y}y\left(c_{y}-y-a_{y}(x-b_{y})^{2}\right),\end{cases}$ (3)
-    where $r_{x}$, $r_{y}$, $a_{x}$, $a_{y}$, $c_{x}$, $c_{y}$ are positive real constants and $b_{x}$, $b_{y}$ are real constants.
 Here, there is the assumption of intraspecific competition, where $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$.
-Moreover, the interaction is mutualistic as long as $y$ remains below the threshold $b_{x}$, but when $y$ exceeds this value, $\frac{\partial f}{\partial y}$ turns negative.
+Moreover, the interaction is mutualistic as long as $y$ remains below the threshold $b_{1}$, but when $y$ exceeds this value, $\frac{\partial f}{\partial y}$ turns negative.
 This dynamic highlights how the strength and nature of the interaction shift depending on species densities.
   • (b)
-    Neuhauser and Fargione, (2004) introduce a model describing facultative mutualism between plants ($x$) and mycorrhizal fungi ($y$).
-In this formulation, the effects on species $y$ remain constant, regardless $x$ or $y$, with intraspecific competition always negative ($\frac{\partial g}{\partial y}<0$) and the mutualistic benefit from $x$ always positive ($\frac{\partial g}{\partial x}>0$).
-Similarly, intraspecific competition for $x$ is constant and negative ($\frac{\partial f}{\partial x}<0$).
+    Neuhauser and Fargione (2004) introduce a model describing facultative mutualism between plants ($x$) and mycorrhizal fungi ($y$).
+In this formulation, the effects on species $y$ remain constant, regardless $x$ or $y$, with intraspecific competition always negative $\left(\frac{\partial g}{\partial y}<0\right)$ and the mutualistic benefit from $x$ always positive $\left(\frac{\partial g}{\partial x}>0\right)$.
+Similarly, intraspecific competition for $x$ is constant and negative $\left(\frac{\partial f}{\partial x}<0\right)$.
 However, the term $\frac{\partial f}{\partial y}$, which characterizes the nature of the interaction, varies depending on the densities of both species.
@@
 $x$ and $y$ must increase together to prevent the relationship from becoming parasitic.
-The system is given by:
-    $\begin{cases}\dot{x}=r_{x}x\left(\frac{K_{x}\gamma_{xy}y-x}{K_{x}+\gamma_{xy}y}-ay\right),\\ \dot{y}=r_{y}y\left(\frac{K_{y}+\alpha_{yx}x-y}{K_{y}}\right),\end{cases}$ (4)
-    where all parameters are positive real constants.
   • (c)
-    The model proposed by Graves et al., (2006) represents obligate mutualism for lichens, where the benefits of the interaction are tightly linked to the densities of both species.
+    The model proposed by Graves et al.
+(2006) represents obligate mutualism for lichens, where the benefits of the interaction are tightly linked to the densities of both species.
 The system assumes constant negative intraspecific competition, with identical functional forms for both species, and the mutualistic interaction varies with the density of the other species.
 Specifically, as the density of $y$ increases, the effect on $x$ decreases, but without any saturation effect—just a continuous increase with density.
-    The system is given by:
-    $\begin{cases}\dot{x}=x\left(r_{x0}+(r_{x1}-r_{x0})\left(1-\exp{(-k_{x}y)}\right)-a_{x}x\right),\\ \dot{y}=y\left(r_{y0}+(r_{y1}-r_{y0})\left(1-\exp{(-k_{y}x)}\right)-a_{y}y\right).\end{cases}$ (5)
-    where $r_{x1}$, $r_{y1}$, $a_{x}$, $a_{y}$, $k_{x}$, $k_{y}$ are positive real constants and $r_{x0}\leq r_{x1}$, $r_{y0}\leq r_{y1}$, $r_{x0}$ and $r_{y0}$ are real constants.
 Here, the interaction benefits are regulated by an exponential term, which prevents unbounded growth as species densities increase.
   • (d)
-    A more general consumer-resource approach is presented by Holland and DeAngelis, (2010), which models the bidirectional exchange of benefits in plant-mycorrhizal systems.
+    A more general consumer-resource approach is presented by Holland and DeAngelis (2009, 2010), which models the bidirectional exchange of benefits in plant-mycorrhizal systems.
 In this framework, one species provides resources while the other enhances nutrient uptake efficiency.
 The system follows a similar functional form for both species, but the dynamics are more complex as all aspects—including the intraspecific competition—vary with the densities of both species.
-    The system is given by:
-    $\begin{cases}\dot{x}=x\left(r_{x}+c_{x}\left(\frac{a_{xy}y}{h_{y}+y}\right)-q_{x}\left(\frac{\beta_{xy}y}{e_{x}+x}\right)-s_{x}x\right),\\ \dot{y}=y\left(r_{y}+c_{y}\left(\frac{a_{yx}x}{h_{x}+x}\right)-q_{y}\left(\frac{\beta_{yx}x}{e_{y}+y}\right)-s_{y}y\right).\end{cases}$ (6)
-    where all parameters are positive constants, except $r_{x}$ and $r_{y}$, which are real constants.
-In this model, the mutualistic interaction is obligate when the intrinsic growth rates are negative ($r_{x}\leq 0$, $r_{y}\leq 0$).
+In this model, the mutualistic interaction is obligate when the intrinsic growth rates are negative ($r_{1}\leq 0$, $r_{2}\leq 0$).
 The interaction structure is dynamic, with the intraspecific competition for $x$ decreasing as $y$ grows, but increasing as $x$ increases.
@@
 White points represent repulsive points, black points represent attractive points and gray points represent saddle points.
-(4(a)) Parameters for Zhang, (2003) model, in (3): $c_{x}=7$, $c_{y}=8$, $a_{x}=0.7$, $a_{y}=0.5$, $b_{x}=4.5$, $b_{y}=3$, $r_{x}=7$, $r_{y}=4$.
-(4(b)) For Neuhauser and Fargione, (2004) model, in (4): $r_{x}=1.4$, $r_{y}=1$, $K_{x}=2$, $K_{y}=5$, $\gamma_{xy}=4$, $\alpha_{xy}=1.5$, $a=0.2$.
-(4(c)) For Graves et al., (2006) model, in (5): $r_{x0}=-2$, $r_{y0}=0.5$, $r_{x1}=4$, $r_{y1}=5$, $k_{x}=0.2$, $k_{y}=0.2$, $a_{x}=0.35$, $a_{y}=0.5$.
-(4(d)) For Holland and DeAngelis, (2010) model, in (6) : $r_{x}=2$, $r_{y}=3$, $c_{x}=3$, $c_{y}=1$, $a_{xy}=1$, $a_{yx}=2.6$, $h_{x}=1$, $h_{y}=1$, $q_{x}=1$, $q_{y}=2.7$, $\beta_{xy}=1$, $\beta_{yx}=1.5$, $e_{x}=0.5$, $e_{y}=3$, $s_{x}=0.7$, $s_{y}=0.5$.
+(4(a)) Parameters for Zhang (2003) model: $c_{1}=7$, $c_{2}=8$, $a_{1}=0.7$, $a_{2}=0.5$, $b_{1}=4.5$, $b_{2}=3$, $r_{1}=7$, $r_{2}=4$.
+(4(b)) For Neuhauser and Fargione (2004) model: $r_{1}=1.4$, $r_{2}=1$, $K_{1}=2$, $K_{2}=5$, $\gamma_{12}=4$, $\alpha_{12}=1.5$, $a=0.2$.
+(4(c)) For Graves et al.
+(2006) model: $r_{10}=-2$, $r_{20}=0.5$, $r_{11}=4$, $r_{21}=5$, $k_{1}=0.2$, $k_{2}=0.2$, $a_{1}=0.35$, $a_{2}=0.5$.
+(4(d)) For Holland and DeAngelis (2010) model: $r_{1}=2$, $r_{2}=3$, $c_{1}=3$, $c_{2}=1$, $a_{12}=1$, $a_{21}=2.6$, $h_{1}=1$, $h_{2}=1$, $q_{1}=1$, $q_{2}=2.7$, $\beta_{12}=1$, $\beta_{21}=1.5$, $e_{1}=0.5$, $e_{2}=3$, $s_{1}=0.7$, $s_{2}=0.5$.
 4 Periodic solutions in the dynamics of two mutualistic species
-A system of the form (1) can represent either predator-prey, competition, or mutualistic interactions. The results described so far do not distinguish between these different types of interactions.
 In this section, we clarify the role of extended mutualism (Definition 2.1) and contrast it with the more constrained definition of strict mutualism.
@@
 Centers are equilibrium points where the trajectories of the system form closed, periodic orbits, but they are neither attractive nor repulsive.
-As a result, the behavior of the system is highly sensitive to initial conditions: each new set of initial conditions produces a distinct trajectory, making the system highly sensitive to small variations (May,, 1972).
+As a result, the behavior of the system is highly sensitive to initial conditions: each new set of initial conditions produces a distinct trajectory, making the system highly sensitive to small variations (May, 1972).
 However, even in the case of centers, when the Lotka-Volterra predator-prey model is modified to represent mutualistic interactions, the centers disappear.
@@
 We identify two different scenarios where the nature of the interaction is not fixed, i.e., where $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ change signs, leading to the emergence of a limit cycle.
+In a related but more specific modeling framework, Song et al.
+(2025) described another example of limit cycle dynamics, based on a detailed mutualism model in which variations in interaction costs and benefits are explicitly analyzed, and shown to induce or suppress cyclic behavior.
 We also introduce the assumption that intraspecific competition can change sign and become intraspecific cooperation (for example, $\frac{\partial g}{\partial y}$ may change sign).
@@
 The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange.
-By assumption 1.9, $(x^{*},y^{*})$ is a repulsive point (the white point), which implies, by Theorem 3.1, that $(x_{2},0)$ is a saddle point (the grey points), as well as $(0,0)$.
-Remark 4.3.
-These assumptions are more restrictive than those in Section 3.1 and imply assumptions H1 to H4.
-Under conditions 1.1-1.10, we have the following result, which will be proven in Section 6.
-Theorem 4.3.
-Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 1.1-1.10. Then a limit cycle exists inside the positive quadrant.
-Example 4.1.
-Let us consider the following system:
-$\left\{\begin{array}[]{ll}\dot{x}&=xf\left(x,y\right)=x\left(a_{1}-b_{1}\left(y-c_{1}\right)^{2}-d_{1}x\right),\\ \dot{y}&=yg\left(x,y\right)=y\left(-a_{2}-b_{2}\left(y-c_{2}\right)^{2}+d_{2}x\right),\end{array}\right$.
-(7)
-with $(a_{i},b_{i},c_{i},d_{i},i\in\{1,2\})$ being positive constants.
-For a certain range of parameters, the functions $f$ and $g$ will satisfy assumptions 1.1 to 1.10.
-However, according to assumptions 1.9 and 1.10, $\Gamma_{f}$ and $\Gamma_{g}$ must intersect while both being decreasing in order to obtain a repulsive equilibrium point.
-Lowering the threshold $c_{2}$, increasing the threshold $c_{1}$, or decreasing the interaction strength $b_{1}$ leads to a shift in the equilibrium location, making it attractive instead.
-System (7) is illustrated in Figure 6.
-Figure 6: Phase portrait leading to a limit cycle.
-Simulation of system (7) with parameters: $a_{1}=3.7$, $b_{1}=0.5$, $c_{1}=2$, $d_{1}=0.6$ and $a_{2}=3$, $b_{2}=0.2$, $c_{2}=4$, $d_{2}=1$.
-The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange.
-The white point represents a repulsive point and gray points represent saddle points.
-The purple trajectory starts at a population density of 2 for species $x$ and 6 for species $y$.
-It converges toward the limit cycle from the outside, moving away from the extinction states, either of both species or the extinction of $y$.
-The gray trajectory starts at densities of 7 for species $x$ and 10 for species $y$, and converges toward the limit cycle from the inside, moving away from the interior equilibrium point.
+By assumption 1.9, $(x^{*},y^{*})$ is a repulsive point (the white point), which implies, by Theorem 3.1, that $(x_{2},0)$ is a saddle point (the grey point), as well as $(0,0)$.
 For our second configuration, we make the following assumptions:
@@
 Remark 4.5.
-As before, these assumptions are more restrictive than those presented in Section 3.1 and imply assumptions H1 to H4.
-Under conditions 2.1-2.10, we have the following result, which will be proven in Section 6.
-Theorem 4.4.
-Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 2.1-2.10. Then a cycle limit exists inside the positive quadrant.
+These assumptions 1.1-1.10 and 2.1-2.9 are more restrictive than those in Section 3.1 and imply assumptions H1 to H4.
+Under conditions 1.1-1.10 or 2.1-2.9, we have the following result, which will be proven in Section 6.
+Proposition 4.6.
+Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 1.1-1.10 or 2.1-2.9.
+Then a limit cycle exists inside the positive quadrant.
+Example 4.7.
+Let us consider the following system:
+$\left\{\begin{array}[]{ll}\dot{x}&=xf\left(x,y\right)=x\left(a_{1}-b_{1}\left(y-c_{1}\right)^{2}-d_{1}x\right),\\ \dot{y}&=yg\left(x,y\right)=y\left(-a_{2}-b_{2}\left(y-c_{2}\right)^{2}+d_{2}x\right),\end{array}\right$.
+(3)
+with $(a_{i},b_{i},c_{i},d_{i},i\in\{1,2\})$ being positive constants.
+For a certain range of parameters, the functions $f$ and $g$ will satisfy assumptions 1.1 to 1.10.
+However, according to assumptions 1.9 and 1.10, $\Gamma_{f}$ and $\Gamma_{g}$ must intersect while both being decreasing in order to obtain a repulsive equilibrium point.
+Lowering the threshold $c_{2}$, increasing the threshold $c_{1}$, or decreasing the interaction strength $b_{1}$ leads to a shift in the equilibrium location, making it attractive instead.
+System (3) is illustrated in Figure 7.
+Figure 7: Phase portrait leading to a limit cycle.
+Simulation of system (3) with parameters: $a_{1}=13.69$, $b_{1}=0.25$, $c_{1}=4$, $d_{1}=0.36$ and $a_{2}=9$, $b_{2}=0.04$, $c_{2}=16$, $d_{2}=1$.
+The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange.
+The white point represents a repulsive point and gray points represent saddle points.
+The purple trajectory starts at a population density of 2 for species $x$ and 6 for species $y$.
+It converges toward the limit cycle from the outside, moving away from the extinction states, either of both species or the extinction of $y$.
+The gray trajectory starts at densities of 7 for species $x$ and 10 for species $y$, and converges toward the limit cycle from the inside, moving away from the interior equilibrium point.
+In both models presented above, Assumption 1.2 (in the first configuration) and Assumption 2.2 (in the second configuration) impose a change of sign in the intraspecific interaction term.
+More precisely, for at least one species, the partial derivative of its per capita growth rate with respect to its own density satisfies in the first configuration $\frac{\partial f}{\partial x}>0$ at low density and $\frac{\partial f}{\partial x}<0$ at high density, and analogously for $g$ in the second configuration.
+Biologically, this corresponds to the presence of an Allee effect: the species experiences positive density dependence at low population size, whereas intraspecific competition dominates at high density.
+This sign change induces a non-monotonic self-regulation mechanism.
+Our analysis suggests that such a mechanism is structurally linked to the existence of a limit cycle.
+In particular, in the absence of any sign change in the intraspecific terms for both species, oscillations cannot occur.
+Proposition 4.8.
+Let a dynamical system on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1).
+Assume that the functions $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$ do not change sign in the positive quadrant.
+Then the system does not admit any limit cycle inside the positive quadrant.
+In particular, a necessary condition for the existence of a limit cycle is that at least one species exhibits a sign change in its intraspecific effect.
+Biologically, this may correspond to an Allee-type positive density dependence at low density.
+The result, proved in Section 6, follows from a variant of the Dulac–Bendixson criterion (see Perko (2013)).
 4.3 From Parasitism to Mutualism: Parameter-Driven Transitions in Interaction Dynamics
-In this section, we include models that, while not always strictly mutualistic, satisfy Definition 2.1, and fit to our general framework.
-Depending on the position in the phase portrait and the population densities of each species, a model satisfying these assumptions could represent a parasitic interaction in one region, followed by a competitive interaction in another region and with different species densities, then a mutualistic one, and then return to a parasitic interaction at yet other densities.
-As Theorem 3.1 is stated in a more general context, it remains valid despite the specific transition dynamics examined in Section 2.2.
-Additionally, Corollaries 3.1 and 3.2 characterize the nature of equilibrium points with index +1+1 when intraspecific competition remains constant in sign.
-However, a given model representing a specific type of interaction can describe an entirely different interaction when its parameter values change (as for example intraspecific competition) while still remaining within the framework of Theorem 3.1.
-In this section, we examine a first model which, depending on the signs of its coefficients—and thus its partial derivatives—can represent either a parasitic or a mutualistic interaction, as discussed in Section 2.2.
+In this section, we illustrate how our framework can be applied to models that cannot be captured by a definition of strict mutualism.
+We propose two models that, while not always strictly mutualistic, satisfy Definition 2.1, and fit to our general framework.
+Depending on population densities and on the location of the equilibrium in the phase portrait, such models may represent parasitic, competitive, or mutualistic interactions in different regions.
+In particular, Definition 2.1 allows for systems in which the nature of species interactions may change continuously as parameters vary, while still remaining within the same mathematical structure.
+Importantly, the general results established in Theorem 3.1 apply independently of transitions between mutualism and parasitism, as the theorem is formulated under assumptions that are broader than those imposed in Section 2.2.
+As a consequence, a model within this framework may describe qualitatively different ecological interactions solely through continuous changes in parameter values—such as the strength or sign of intraspecific competition—without leaving the scope of Theorem 3.1.
+First, we examine a model which, depending on the signs of its coefficients—and thus its partial derivatives—can represent either a parasitic or a mutualistic interaction, as discussed in Section 2.2.
 We analyze the consequences of this change on phase portraits and equilibrium points of index +1+1.
@@
 Our first example is the following system:
-$\begin{cases}\dot{x}=\delta x\left(a_{x}-b_{x}\left(x-c_{x}\right)^{2}-d_{x}y\right),\\ \dot{y}=\delta y\left(a_{y}-b_{y}\left(y-c_{y}\right)^{4}-d_{y}x\right),\end{cases}$ (8)
+$\begin{cases}\dot{x}=\delta x\left(a_{1}-b_{1}\left(x-c_{1}\right)^{2}-d_{1}y\right),\\ \dot{y}=\delta y\left(a_{2}-b_{2}\left(y-c_{2}\right)^{4}-d_{2}x\right),\end{cases}$ (4)
 where:
@@
 White points represent repulsive points, black points represent attractive points and gray points represent saddle points.
-With: $a_{x}=8$, $b_{x}=1$, $c_{x}=5.5$, $d_{x}=1$ and $a_{y}=9.5$, $b_{y}=0.5$, $c_{y}=4$, $d_{y}=1$.
+With: $a_{1}=8$, $b_{1}=1$, $c_{1}=5.5$, $d_{1}=1$ and $a_{2}=9.5$, $b_{2}=0.5$, $c_{2}=4$, $d_{2}=1$.
 In our second example, we do not change the sign of the partial derivatives; the model has two regions, a mutualistic region and a parasitic region.
@@
 We use the following model:
-$\begin{cases}\dot{x}=x(-a_{x}-b_{x}(x-c_{x})^{2}+d_{x}y),\\ \dot{y}=y(a_{y}-b_{y}(x-c_{y})^{4}-d_{y}y).\end{cases}$ (9)
+$\begin{cases}\dot{x}=x(-a_{1}-b_{1}(x-c_{1})^{2}+d_{1}y),\\ \dot{y}=y(a_{2}-b_{2}(x-c_{2})^{4}-d_{2}y).\end{cases}$ (5)
 Figure 9 provides two phase portraits of the same model, demonstrating how varying parameter values affect equilibrium locations.
@@
 In contrast, in Figure 9(b), stable coexistence occurs within a region of strict mutualism.
-(a) Phase portrait of (9), with parameters: $a_{x}=1.7$, $c_{x}=5.2$ and $c_{y}=5$.
-(b) Phase portrait of (9), with parameters: $a_{x}=1.5$, $c_{x}=1.9$ and $c_{y}=7$.
+(a) Phase portrait of (5), with parameters: $a_{1}=1.7$, $c_{1}=5.2$ and $c_{2}=5$.
+(b) Phase portrait of (5), with parameters: $a_{1}=1.5$, $c_{1}=1.9$ and $c_{2}=7$.
 Figure 9: Phase portraits illustrating how changing parameter values—while maintaining the interaction type—modifies the equilibrium locations.
@@
 Black points represent attractive points and gray points represent saddle points.
-System (9) with: $b_{x}=0.4$, $d_{x}=1$, $a_{y}=6.5$, $b_{y}=0.03$ and $d_{y}=1$.
+System (5) with: $b_{1}=0.4$, $d_{1}=1$, $a_{2}=6.5$, $b_{2}=0.03$ and $d_{2}=1$.
 5 Discussion
 In this study we investigated the deterministic dynamics of two interacting species under density-dependent interactions, with a particular focus on mutualistic interactions.
-Our approach allowed to draw a general theoretical framework, formalized by Theorem 3.1, which is independent of the specific type of ecological interaction considered, covering for instance predator-prey, competitive and mutualistic interactions.
-However, while most mutualistic models reviewed in Hale and Valdovinos, (2021) satisfy assumptions H1 to H4, some mutualistic models do not necessarily adhere to these assumptions.
-For instance, they may involve more than two isoclines intersecting at a single point, or include emigration effects (Thompson et al.,, 2006; Holland and DeAngelis,, 2010).
-Our framework nevertheless provides general predictions of the dynamics of species involved in a wide array of ecological interactions, and highlights the key-role of density-dependent processes, that are likely widespread in natural communities (Kawatsu and Kondoh,, 2018).
-We extended the classical definition of mutualism, allowing density-dependent transitions between mutualistic and antagonistic interactions.
+Our approach allowed to draw a general theoretical framework, formalized by Theorem 3.1, which is independent of the specific type of ecological interaction considered, covering for instance predator-prey, competitive and a wide range of mutualistic interactions, as highlighted in Tables 1, 2 and 3.
+However, while most mutualistic models reviewed in Hale and Valdovinos (2021) satisfy assumptions H1 to H4, some mutualistic models do not necessarily adhere to these assumptions: for instance, they may involve more than two isoclines intersecting at a single point, or include emigration (Thompson et al., 2006; Holland and DeAngelis, 2010).
+Importantly, we extended the classical definition of mutualism, allowing density-dependent transitions between mutualistic and antagonistic interactions.
+Such transition in the type of ecological interactions are well documented at the evolutionary scale (see for instance Drew et al.
+(2021)), and could be favored in ecological context where relative densities among interacting species might change.
 We thus proposed a broader Definition 2.1: a system is considered mutualistic if there exists at least one region for each species in the phase portrait where each benefits from the presence of the other.
-This relaxed definition enables the exploration of more diverse ecological dynamics and highlights biologically realistic features of species interactions.
-In textbook example of mutualistic interactions, such as pollination of flowering plants by insects, large densities of invasive bees have been documented to decrease plant reproductive success because of the damages caused by over-visitation of flowers (Aizen et al.,, 2014).
+This relaxed definition enables the exploration of more diverse ecological dynamics and highlights biologically realistic features of species interactions, that can crucially depend on the relative densities of interacting species (see for instance density-dependent mutualistic effects in ant-aphid interactions, Breton and Addicott (1992)).
+Even in the textbook examples of mutualistic interactions, such as pollination of flowering plants by insects, large densities of invasive bees have been documented to decrease plant reproductive success because of the damages caused by over-visitation of flowers (Aizen et al., 2014).
 As shown in some of the mutualistic models we studied, it is possible that small changes in parameter values could cause an equilibrium point that was initially located in a strictly mutualistic region to shift into a parasitic one.
-By allowing such transitions between mutualism and parasitism, our extended framework opened to the analysis of more complex systems and revealed new dynamical phenomena such as the emergence of limit cycles.
-Our theoretical predictions therefore provide a more realistic view of population dynamics of interacting species, opening new research avenues for community ecology.
+By allowing such transitions between mutualism and parasitism, our extended framework sheds new light on the qualitative structure of species interactions and permits the emergence of dynamical phenomena not accessible under strictly mutualistic assumptions, including the theoretical possibility of limit cycles.
+Numerous empirical studies report cyclic population dynamics in nature, which are overwhelmingly interpreted as the outcome of predator–prey or parasitic interactions (Turchin, 2003, p.
+383-384).
+We showed that cycles can indeed not occur in case of strict mutualism.
+However, we also showed that periodic oscillations can occur when there are transitions between mutualistic and parasitic interaction regimes.
+To our knowledge, three contexts can give rise to periodic oscillations in a mutualism model: the two general contexts we give in Section 4.2 as well as the model analyzed by Song et al.
+(2025).
+Note that Song et al.
+(2025)’s model is a particular case of System 1 but it does not enter into the general context of Section 4.2, thus suggesting that periodic oscillations can occur in other situations.
+It is striking however that our examples and Song et al.
+(2025)’s model share a common feature: an Allee effect, i.e.
+a positive intraspecific feedback interactions when density is low.
+This observation raises the question of whether, beyond mutualism–parasitism transitions, an Allee effect may also be a necessary condition for the emergence of such cycles.
+Overall, our results propose an alternative and complementary interpretative framework for explaining some of the cyclic patterns observed in natural systems.
+To our knowledge there is no empirical evidence of mutualistic interactions exhibiting sustained population cycles, but long-term survey of mutualistic populations with precise estimations of densities are actually scarce.
+Given the current lack of precise empirical data on the temporal dynamics of mutualistic populations, our results should primarily be interpreted as theoretical possibilities, which might nonetheless become relevant under particular ecological or evolutionary scenarios yet to be documented.
+Our framework thus highlights potential behaviors that could guide future empirical investigations aimed at identifying conditions under which mutualistic interactions may fluctuate rather than converge to equilibrium.
+Finally, our framework provides general predictions of the dynamics of species involved in a wide array of ecological interactions, and highlights the key-role of density-dependent processes, that are widespread in natural communities (Kawatsu and Kondoh, 2018).
+In particular, the framework shows how density dependence can shape interaction outcomes across a continuum of ecological contexts.
+It does not rely on specific algebraic expressions for interaction terms but only on qualitative assumptions on the signs of partial derivatives, allowing it to encompass a broad family of models without requiring explicit functional forms.
+This flexibility opens the way to future extensions incorporating environmental variability, state-dependent interactions, and slow evolutionary changes in parameters.
 6 Proofs
-In this section, we provide the proofs of the different results stated in Section 3 and 4.
-We begin with the proof of Theorem 3.1, which describes the alternation of equilibrium point indices along the isoclines.
+In this section, we provide the proofs of the different results stated in Sections 3 and 4.
+We begin with the proof of Theorem 3.1 and Proposition 3.5, which describe the alternation of equilibrium point indices along the isoclines.
 Proof of Theorem 3.1.
@@
 ∎
+Proof of Proposition 3.5.
+We restrict the analysis to the case where the equilibrium point lies on the $x$-axis i.e.
+is generated by an intersection of the isocline $\Gamma_{f}$ with the axis $y=0$.
+The case of an equilibrium point located on the $y$-axis, arising from an intersection of $\Gamma_{g}$ with the axis $x=0$, follows by symmetry.
+Let $(x_{1},0)$ denote the equilibrium point closest to the origin $(0,0)$.
+Under our assumptions, there may exist at most one additional equilibrium point on the $x$-axis, denoted by $(x_{2},0)$, with $x_{1}<x_{2}$, which may or may not exist.
+We keep the same notations and conventions as in the proof of Theorem 3.1.
+In a neighborhood of the equilibrium point $(x_{1},0)$, four local configurations of the vector field may occur, depending on the associated sign of the vector field across the adjacent regions.
+The equilibrium point $(x_{1},0)$ may have index +1+1, corresponding to an attractive or repulsive equilibrium.
+In this case, the alternation of signs across the four regions is either (+,),(,)(+,-),(-,-) or (,+),(+,+)(-,+),(+,+) illustrated in Figure 10(10(a)).
+Alternatively, $(x_{1},0)$ may have index 1-1, in which case it is a saddle point, and the alternation of signs is either (+,+),(,+)(+,+),(-,+) or (,),(+,)(-,-),(+,-) illustrated in Figure 10(10(b)).
+As in the interior case, the ordering of signs is read clockwise.
+(a) Local possible structure of the phase portrait around an index +1+1 equilibrium on the $x$-axis.
+(b) Local possible structure of the phase portrait around an index 1-1 equilibrium on the $x$-axis.
+Figure 10: Four possible local configurations of the phase portrait near an equilibrium point located on the $x$-axis., Two possible configurations are shown for an equilibrium point with index +1+1 (10(a)) and index 1-1 (10(b)).
+The alternation of signs of the vector field across adjacent regions is read clockwise.
+These configurations correspond to equilibrium points generated by intersections of the isocline $\Gamma_{f}$ with the axis $x=0$.
+Thus, when following the isocline $\Gamma_{f}$ and encountering an equilibrium point located in the strictly positive quadrant, the first intersection of $\Gamma_{f}$ with $\Gamma_{g}$ reached when moving along $\Gamma_{f}$ starting from $(x_{1},0)$, the alternation of signs is reversed, while still being read clockwise.
+More precisely, if the equilibrium point $(x_{1},0)$ on the $x$-axis is characterized by the alternation (+,),(,)(+,-),(-,-) in Figure 10(a), then the first equilibrium point encountered in the strictly positive quadrant exhibits the alternation $(+,+)$, (,+)(-,+), (,)(-,-), and (+,)(+,-), corresponding to case $(ii)$ in the proof of Theorem 3.1, which yields an equilibrium point with index 1-1.
+The remaining cases follow the same principle.
+The remaining alternations in the strictly positive quadrant then follow from Theorem 3.1.
+It remains to determine the nature of the equilibrium point $(x_{2},0)$, when it exists.
+At this stage, we have shown that the index alternation persists when the equilibrium point $(x_{1},0)$ on the $x$-axis is taken into account.
+It remains to determine whether this alternation is also satisfied by a second equilibrium point on the axis, if such a point exists.
+This depends on whether the isocline $\Gamma_{g}$ intersects the $x$-axis, and on the number of such intersections.
+Indeed, even if intersections of $\Gamma_{g}$ with the $x$-axis do not generate additional equilibrium points, they modify the sign of the vector field in the vertical direction (for example changing a $(+,+)$ into a (+,)(+,-)).
+If $\Gamma_{g}$ does not intersect the $x$-axis between $(x_{1},0)$ and $(x_{2},0)$, or if $\Gamma_{g}$ intersects it twice on this interval (resulting in a double sign change), then the alternation at $(x_{2},0)$ is the reverse of that at $(x_{1},0)$.
+For instance, if the alternation at $(x_{1},0)$ is $(+,+)$ and (,+)(-,+) in Figure 10(b), then at $(x_{2},0)$ it becomes (,+)(-,+) and $(+,+)$, which transforms an equilibrium point with index 1-1 into one with index +1+1.
+It remains to verify whether this behavior is consistent with the previously established index alternation.
+In both situations (zero or two intersections of $\Gamma_{g}$ with $x$-axis), $\Gamma_{g}$ may generate equilibrium points in the phase portrait through its intersections with $\Gamma_{f}$.
+Since the isoclines are assumed to be continuous and by Theorem 3.1, the number of equilibrium points created by intersections of $\Gamma_{f}$ and $\Gamma_{g}$ before $\Gamma_{f}$ reaches $(x_{2},0)$ is necessarily even.
+Consequently, the last equilibrium point encountered in the strictly positive quadrant, corresponding to the final intersection of $\Gamma_{f}$ and $\Gamma_{g}$, has the same index as $(x_{1},0)$.
+As shown above, the equilibrium point $(x_{2},0)$ has an index opposite to that of $(x_{1},0)$.
+Therefore, the alternation of equilibrium point indices is preserved when $\Gamma_{g}$ has zero or two intersections with the $x$-axis between $(x_{1},0)$ and $(x_{2},0)$.
+It remains to consider the case where the isocline $\Gamma_{g}$ intersects the $x$-axis exactly once between $(x_{1},0)$ and $(x_{2},0)$.
+By the same reasoning as above, for an equilibrium point $(x_{1},0)$ characterized by a given alternation of sign pairs, the intersection of $\Gamma_{g}$ with the $x$-axis induces a change in the sign of the vector field in the $y$-direction along the axis.
+Then, when reaching $(x_{2},0)$, the intersection with $\Gamma_{f}$ induces a change in the sign of the vector field in the $x$-direction.
+For instance, if the alternation around $(x_{1},0)$ is given by $(+,+)$ and (,+)(-,+), then after crossing $\Gamma_{g}$ the alternation becomes (,)(-,-), and after crossing $\Gamma_{f}$ it becomes (+,)(+,-) at $(x_{2},0)$.
+As a result, the equilibrium point $(x_{2},0)$ has the same index as $(x_{1},0)$.
+Moreover, since $\Gamma_{g}$ is continuous, it intersects the isocline $\Gamma_{f}$ at least once between $(x_{1},0)$ and $(x_{2},0)$, and in fact an odd number of times.
+Consequently, the last equilibrium point generated by an intersection of $\Gamma_{f}$ and $\Gamma_{g}$ has an index opposite to that of $(x_{1},0)$, and hence opposite to that of $(x_{2},0)$.
+Therefore, the alternation of equilibrium point indices is preserved when following $\Gamma_{f}$ from $(x_{1},0)$ to $(x_{2},0)$.
 We now provide the proof of Corollary 3.6, which is a consequence of Theorem 3.1. The proof of Corollary 3.9 is very similar and will be omitted.
 Proof of Corollary 3.6.
-Let $\left(x,y\right)$ an equilibrium point in the strictly positive quadrant. We can calculate the Jacobian of the system (1) at the point $\left(x,y\right)$:
+Let $\left(x,y\right)$ be an equilibrium point in the strictly positive quadrant. We can calculate the Jacobian of the system (1) at the point $\left(x,y\right)$:
 $J\left(x,y\right)=\begin{pmatrix}x\frac{\partial f\left(x,y\right)}{\partial x}&x\frac{\partial f\left(x,y\right)}{\partial y}\\ y\frac{\partial g\left(x,y\right)}{\partial x}&y\frac{\partial g\left(x,y\right)}{\partial y}\end{pmatrix}$.
@@
 It implies that the point $\left(x,y\right)$ is an attractive node if $\text{det}\left(J\left(x,y\right)\right)>0$, else, if $\text{det}\left(J\left(x,y\right)\right)<0$, the point $\left(x,y\right)$ is a saddle point.
+∎
+We now prove Corollary 3.8. Notice that the proof of Corollary 3.10 will be omitted since it follows exactly the same arguments.
+Proof.
+We suppose that $x_{1}>0$ and $(x_{1},0)$ is an equilibrium point on the $x$-axis, that is, $f(x_{1},0)=0$.
+We prove the result for equilibria on the $x$-axis, i.e.
+for equilibrium points given by the intersection of $\Gamma_{f}$ with the line $y=0$.
+The proof for equilibria on the $y$-axis, corresponding to the intersection of $\Gamma_{g}$ with the line $x=0$, is identical.
+We recall that $f\in\operatorname{\mathcal{C}^{1}}$ and that we assume $\frac{\partial f}{\partial x}(x,y)<0\quad\text{for all }(x,y)$.
+We assume by contradiction that there exist two distinct equilibria $0<x_{1}<x_{2}$ on the $x$-axis such that $f(x_{1},0)=f(x_{2},0)=0$.
+Since $f$ is $C^{1}$, for $h>0$ a first-order Taylor expansion of $f(\cdot,0)$ at $x_{1}$ yields
+$f(x_{1}+h,0)=f(x_{1},0)+h\frac{\partial f}{\partial x}(x_{1},0)+o\mathopen{}\left(h\right)=h\frac{\partial f}{\partial x}(x_{1},0)+o\mathopen{}\left(h\right),$
+and
+$f(x_{1}-h,0)=-h\frac{\partial f}{\partial x}(x_{1},0)+o\mathopen{}\left(h\right).$
+Since $\frac{\partial f}{\partial x}(x_{1},0)<0$, we have $f(x,0)<0$ for $x>x_{1}$ sufficiently close to $x_{1}$, and $f(x,0)>0$ for $x<x_{1}$ sufficiently close to $x_{1}$.
+Similarly, a Taylor expansion at $x_{2}$ implies that $f(x,0)>0$ for $x<x_{2}$ sufficiently close to $x_{2}$.
+By continuity of $f(\cdot,0)$, there must then exist $x\in(x_{1},x_{2})$ such that $f(x,0)=0$, which contradicts hypotheses H1 to H4, which allow at most two equilibrium points on the axes.
+Hence, the equilibrium on the $x$-axis is unique.
+Let $x_{1}$ be the unique equilibrium on the $x$-axis.
+From the previous sign analysis, $f(x,0)>0$ for $x<x_{1}$ close to $x_{1}$, and $f(x,0)<0$ for $x>x_{1}$ close to $x_{1}$.
+By hypothesis, the index of $(x_{1},0)$ is +1+1, and therefore, by Proposition 3.5, $(x_{1},0)$ is an attractive equilibrium point.
 ∎
@@
 Since $h$ is $\operatorname{\mathcal{C}^{1}}$ and has the same derivative as in the previous proof, we arrive at the same contradiction.
-We end with the proof of Theorem 4.3 and 4.4 which provide conditions under which stable limit cycles may be observed.
-Proof of Theorem 4.3.
-We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 5. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
+We present the proof of Proposition 4.6, which provides conditions under which stable limit cycles may arise.
+Proof of Proposition 4.6.
+The proof is based on the Poincaré-Bendixson theorem (Ciesielski, 2012), used here through the fivefold way formulation introduced by Coleman (1983).
+Our strategy consists in constructing a positively invariant region of the phase plane and analyzing the possible $\omega$-limit sets of a trajectory entering this region.
+For the first case, we construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 5.
+We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
 AA lies along a trajectory, defined by $x=0$, $y>0$, and $(0,0)$ is an equilibrium point.
@@
 Our goal now is to demonstrate that the limit set $\omega(\gamma)$ of the trajectory $\gamma$ is a limit cycle.
-According to the fivefold way (see Coleman, (1983), page 265), this will follow if we can show that $\omega(\gamma)$ contains no critical points.
+According to the Poincaré-Bendixson theorem, stated in the form of the fivefold way (see Coleman (1983), page 265), this will follow if we can show that $\omega(\gamma)$ contains no critical points.
 The critical points of system (1) are located at the origin $(0,0)$, at the point $(x_{2},0)$ on the $x$-axis, and at the point $\left(x^{*},y^{*}\right)$.
-$\left(x^{*},y^{*}\right)$ cannot be in $\omega(\gamma)$ because $\left(x^{*},y^{*}\right)$ is repusilve by assumption 1.9.
+$\left(x^{*},y^{*}\right)$ cannot be in $\omega(\gamma)$ because $\left(x^{*},y^{*}\right)$ is repulsive by assumption 1.9.
 Moreover we cannot have that $\omega(\gamma)=\{(0,0)\}$.
@@
 Thus $\omega(\gamma)$ contains no critical points.
-By the fivefold way (see Coleman, (1983), page 265) we have that $\omega(\gamma)$ is a cycle.
-∎
-Proof of Theorem 4.4.
-We will follow the same approach as in the previous proof.
+By the fivefold way (see Coleman (1983), page 265) we have that $\omega(\gamma)$ is a cycle.
+By the Poincaré-Bendixson theorem, applied through the fivefold way (see Coleman (1983), page 265) we have that $\omega(\gamma)$ is a cycle.
+We will follow the same approach for the second case.
 We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 6.
@@
 ∎
+We conclude with the proof of Proposition 4.8, which provides a necessary condition for the existence of a limit cycle.
+Proof.
+We consider system (1) on the strictly positive quadrant:
+$\left\{\begin{array}[]{ll}\dot{x}=xf(x,y),\\ \dot{y}=yg(x,y).\end{array}\right$.
+We assume that $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$ do not change sign.
+We consider the Dulac function
+$B(x,y)=\frac{1}{xy},$
+which is $C^{1}$ on the strictly positive quadrant.
+Let F(x,y)=(xf(x,y),yg(x,y))F(x,y)=(xf(x,y),\,yg(x,y)) denote the vector field. Then
+$BF=\left(\frac{f(x,y)}{y},\,\frac{g(x,y)}{x}\right).$
+Its divergence is
+$\nabla\cdot(BF)=\frac{\partial}{\partial x}\left(\frac{f(x,y)}{y}\right)+\frac{\partial}{\partial y}\left(\frac{g(x,y)}{x}\right)=\frac{1}{y}\frac{\partial f}{\partial x}+\frac{1}{x}\frac{\partial g}{\partial y}.$
+Since $x>0$ and $y>0$, the sign of $\nabla\cdot(BF)$ is entirely determined by the signs of $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$.
+If both $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$ are negative everywhere (or both postive everywhere), then $\nabla\cdot(BF)$ does not change sign and is not identically zero.
+By the Bendixson–Dulac criterion (see Perko (2013)), a $C^{1}$ planar vector field defined on on the strictly positive quadrant cannot admit a periodic orbit entirely contained in that domain if there exists a $C^{1}$ Dulac function whose weighted divergence does not change sign.
+Therefore, the system admits no limit cycle in the strictly positive quadrant. ∎
 Acknowledgments
@@
 Appendix C Isocline behaviors and Point classifications
-[…]
+We study system (1).
+We assume that the functions $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$, so the sign of their partial derivatives remains strictly constant in a neighborhood of the equilibrium points.
+By the implicit function theorem, we can determine whether the isoclines are increasing or decreasing near these equilibrium points, in addition to identifying their nature using the trace-determinant diagram.
+We classify the results based on the signs of the partial derivatives in Table 4 and specify when the system satisfies the assumptions of strict mutualism as defined in Section 2.2.
 $\frac{\partial g}{\partial x}>0$ $\frac{\partial g}{\partial x}>0$ $\frac{\partial g}{\partial x}<0$ $\frac{\partial g}{\partial x}<0$